### v1

# Compressed Momentum-based Single-Point Zero-Order Algorithm for Stochastic Distributed Nonconvex Optimization

Abstract:
This paper studies a compressed momentum-based single-point zeroth-order algorithm for stochastic distributed nonconvex optimization, aiming to alleviate communication overhead and address the unavailability of explicit gradient information. In the developed framework, each agent has access only to stochastic zeroth-order information of its local objective function, performs local stochastic updates with momentum, and exchanges compressed updates with its neighbors. We theoretically prove that the proposed algorithm achieves the convergence rate $\mathcal{O}(\frac{1}{\sqrt[4]{T}})$ to the exact solution with diminishing step sizes and can achieve a sublinear convergence rate ${\mathcal{O}}(\frac{1}{{\sqrt{T}}})$ towards a neighborhood of the stationary point with fixed step sizes. Numerical experiments validate the effectiveness and communication efficiency of the proposed algorithm.
keywords:
Compressed communication, distributed nonconvex optimization, momentum-based method, zeroth-order algorithm
1 Introduction
With the rapid growth of datasets and the expansion of network scales, distributed optimization has attracted significant attention due to its enhanced parallel computational capabilities and inherent robustness, where agents try to cooperatively solve a problem with information exchange only limited to immediate neighbors in the network. In many networked systems, such as power systems (mao2019finite) and wireless sensor networks (kuthadi2022optimized, lei2022distributed), each computational node is typically limited to accessing information from its immediate neighbors, driven by requirements for security and scalability.
To date, a wide variety of distributed algorithms have been proposed, including distributed gradient tracking algorithms (lei2022distributed, chen2024compressed), distributed ADMM (carnevale2025admm, gholami2023admm), distributed pull-push gradient-based algorithm (nedic2025ab, dimlioglu2025communication) and so on. Additionally, a popular SGD heuristic that has proven to be crucial in many applications is the use of momentum, i.e., the use of a weighted average of past gradients instead of the current gradient (sutskever2013importance). Subsequently, xia2024momentum investigated a paradigm of momentum-based systems for nonconvex constrained optimization and distributed nonconvex optimization. li2025stochastic proposed a gradient tracking stochastic distributed optimization algorithm with adaptive momentum.
However, in fields such as biochemistry, commercial applications and complex engineering systems, accessing gradients may pose significant challenges due to data privacy concerns, black-box constraints, or the high computational costs involved. To handle this problem, the gradient is approximated through random sampling and finite differences in zeroth-order (ZO) optimization. Existing ZO optimization algorithms can be divided into the following categories, namely, ZO with one-point estimator and ZO with two- and multiple-point estimator. flaxman2004online proposed the ZO algorithm with one-point estimator that queries one function value at each iteration to estimate the gradient. Nowadays, some research findings are available to integrate ZO optimization with momentum-based algorithms. For instance, huang2022accelerated studied a class of accelerated zeroth-order and first-order momentum methods for both nonconvex mini-optimization and minimax-optimization. qian2023zeroth explored a ZO proximal stochastic recursive momentum algorithm.
Note that the distributed optimization algorithms require each agent to exchange information with its neighboring agents to obtain the global information. However, the communication channel often has limited bandwidth. Hence, a series of compression techniques have been developed to reduce communication overhead, including standard uniform quantizers (xu2024quantized), unbiased compressors (condat2022ef), and sparsification methods (wangni2018gradient).
In this paper, motivated by the well-known momentum-based optimization algorithm and by the practical limitation that querying multiple function values is often infeasible, especially in stochastic environments, we develop a Compressed Momentum-based Single-point Zeroth-Order (CMSGD_\_1P) algorithm for stochastic distributed nonconvex optimization. The contributions are summarized as follows.
  • We propose a momentum-based method for stochastic distributed nonconvex optimization. Unlike singh2021squarm, we consider a more challenging gradient-free distributed setting, where gradients are approximated via noisy single-point function evaluations. Compared with mhanna2023single, our method incorporates compressed communication, which substantially reduces the communication overhead in distributed networks.
  • We establish that the proposed CMSGD_\_1P algorithm can achieve a convergence rate $\mathcal{O}(\frac{1}{\sqrt[4]{T}})$ towards the exact solution with diminishing step sizes and can achieve the sublinear convergence rate ${\mathcal{O}}(\frac{1}{{\sqrt{T}}})$ towards a neighborhood of a stationary point with fixed step sizes, where $T$ is the number of iterations.
  • In contrast to mhanna2023single, which assumes that $\|x\|$ is uniformly bounded. In this paper, we analyze two related auxiliary components to avoid the need for uniform boundedness assumption on $\|x\|$.
NotationsNotations: Let $\mathbb{R}$ denote the set of real numbers. Vectors with all entries equal to 11 or 0 are denoted by $\mathbf{1}$ and $\mathbf{0}$, respectively. $\|\cdot\|$ and $\|\cdot\|_{F}$ symbolize the $\ell_{2}$–norm and the Frobenius norm, respectively. The set of $n$-dimensional real vectors is denoted by $\mathbb{R}^{n}$, and $\mathbf{I}$ means the identity matrix. For a differentiable function $f$, $\nabla f$ represents its gradient. For two vectors $\in\mathbb{R}^{n}, a,b\langle a,b\rangle$ is the standard inner product. For a positive semi-definite matrix $\mathbf{W}$, $\lambda(\mathbf{W})$ denotes its spectral radius. Additionally, $\operatorname{sgn}(\cdot)$, $\circ$, $\lfloor\cdot\rfloor$, and $|\cdot|$ denote the element–wise sign, Hadamard product, floor function, and absolute value, respectively.
2 Preliminaries and problem formulation
2.1 Problem formulation
Consider a set of agents ${\mathcal{V}}=\{1,2,...,n\}$ connected by a communication network. Each agent $i$ is associated with a local objective function $F_{i}:{\mathbb{R}^{d}}\mapsto\mathbb{R}$. All agents aim to collaboratively find the decision variable $x\in{\mathbb{R}^{d}}$} that minimizes the stochastic global objective function:
$\mathop{\min}\limits_{x\in{\mathbb{R}^{d}}}{f}(x):=\frac{1}{n}\sum\limits_{i=1}^{n}{{F_{i}}(x)},$ (1)
where
${F_{i}}(x)\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}{\mathbb{E}_{\xi_{i}\sim{\mathcal{D}}_{i}}}[{f_{i}}(x,{\xi_{i}})],$ (2)
with $x$ denotes the global decision vector and $\xi_{i}$} is the local data that follows the local distribution ${\mathcal{D}}_{i}$, and ${f_{i}}(x,{\xi_{i}})$ is the stochastic local cost function of each agent ii. In addition, we assume that $\mathbb{E}_{\xi_{i}\sim{\mathcal{D}}_{i}}{{{[{{f_{i}}({x},{\xi_{i}})}]}^{2}}}\leq\gamma_{1}^{2}$, where $\gamma_{1}>0$.
Assumption 1 (The local functions)
We assume the Lipschitz continuity of the local objective functions $f_{i}(x,\xi_{i})$ such that
$\|f_{i}(x,\xi_{i})-f_{i}(x^{\prime},\xi_{i})\|\leq L_{f_{1}}\|x-x^{\prime}\|,\quad\forall x,x^{\prime}\in\mathbb{R}^{d},$ (3)
where $L_{f_{1}}$ denotes the Lipschitz constant.
Assumption 2 (The objective function)
Both $\nabla F_{i}(x)$ and $\nabla^{2}F_{i}(x)$ exist and are continuous, and there exists a constant $L_{{f_{2}}}>0$ such that $\|\nabla^{2}F_{i}(x)\|\leq L_{{f_{2}}}$, $\forall x\in\mathbb{R}^{d}.$ Hence, we can say that the stochastic objective function $f(x)$ is $L_{{f_{2}}}$–smooth:
$f(x)\leq f(x^{\prime})+\langle\nabla f(x^{\prime}),x-x^{\prime}\rangle+\frac{L_{{f_{2}}}}{2}\|x-x^{\prime}\|^{2}.$ (4)
To address the problem (1), we focus on the following aspects: exchanging information with other agents and collecting information about the local cost function. We introduce the following mechanisms to address these two aspects, together with several additional assumptions.
2.2 Graph Theory
An undirected graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ is introduced to describe the communication network with $n$ agents, where $\mathcal{V}=\{1,...,n\}$ is the set of agents. Define ${\mathcal{E}}\subset{\mathcal{V}}\times{\mathcal{V}}$ as the set of communication channels. The set of neighbors of the agent is denoted by ${{\mathcal{N}}_{i}}=\{j|(i,j)\in{\mathcal{E}}\}$. The non-negative weighted matrix $\mathbf{W}$ is represented as $\mathbf{W}=[{w_{ij}}]\in{{\mathbb{R}}^{n\times n}}$. And all diagonal elements $w_{ii}$ are strictly positive. An channel $(i,j)\in{\mathcal{E}}$ indicates that the agent $i$ can communicate information with the agent $j$. Furthermore, ${w_{ij}}>0$ if $(i,j)\in{\mathcal{E}}$ and otherwise ${w_{ij}}=0$.
Assumption 3
Assume that the graph $\mathcal{G}$} is connected. The weighted matrix $\mathbf{W}$ is symmetric and doubly stochastic, i.e., ${\mathbf{1}^{T}}\mathbf{W}={\mathbf{1}^{T}},\mathbf{W}\mathbf{1}=\mathbf{1}$.
As shown in Assumption 3, for such a $\mathbf{W}$ associated with $\mathcal{G}$}, it is known that $\lambda_{1}(\mathbf{W})=1$ and $\lambda_{i}(\mathbf{W})\in(-1,1)$ for all $i=2,\dots,n$. The spectral gap $\delta\in(0,1]$ is defined as $\delta:=1-|\lambda_{2}(\mathbf{W})|$.
2.3 Zeroth-order optimization
This section concentrates on utilizing the stochastic zeroth-order (ZO) information to estimate the gradients of local objective functions. In the absence of explicit gradient expressions, we assume each agent could query the local function value at one point.
Let $F(x)=\mathbb{E}_{\xi}[f(x,\xi)]$, where $f(x,\xi)$ denotes a stochastic differentiable function. As shown in flaxman2004online, the one-point estimator can be constructed as
$g(x,\xi):=\frac{d}{\gamma}f(x+\gamma u,\xi)u,$ (5)
where $x$ means the optimization variable, $\gamma>0$ denotes the parameter that sets the level of smoothing. And $u$ is a random vector that follows a symmetric distribution, $d$ means the problem dimension.
To solve the problem (1), agent $i$ can only query the function values of $f_{i}$ at exactly one point. These function queries are assumed as $\tilde{f}_{i}=f_{i}+\varphi_{i}$. We note that $\varphi_{i}$ is added to differentiate between the stochastic nature of $f_{i}$ and the noise imposed on its querying.
Then, we assume that the function queries are
$\tilde{f}_{i}({x_{i,t}}+{\gamma_{g}}{u_{i,t}},{\xi_{i,t}})=f_{i}({x_{i,t}}+{\gamma_{g}}{u_{i,t}},{\xi_{i,t}})+\varphi_{i,t},$
where $x_{i,t}\in\mathbb{R}^{d}$ is the estimation of $x$ maintained by agent $i$ at time step $t$; $\varphi_{i,t}$ denotes the additive noise; $\gamma_{g}$ is a positive parameter called the smoothing radius; $u_{i,t}$ is a random perturbation vector generated by the agent ii.
Inspired by (5), we have
$g_{i,t}$ $=\frac{d}{\gamma_{g}}u_{i,t}\tilde{f}_{i}\big(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t}\big)$ (6)
$=\frac{d}{\gamma_{g}}u_{i,t}(f_{i}(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t})+\varphi_{i,t}).$
For the noise term $\varphi_{i,t}$ and the perturbation vector $u_{i,t}$, we adopt the following assumptions.
Assumption 4
The noise term $\varphi_{i,t}$ is assumed to have a zero mean and be uncorrelated with bounded variance, where $\mathbb{E}(\varphi_{i,t})=0$ and $\mathbb{E}(\varphi_{i,t}^{2})=\vartheta_{2}<\infty,i\in\mathcal{N}$, and $\mathbb{E}(\varphi_{i,t}\varphi_{j,t})=0$ for $i\neq j$.
Assumption 5
The perturbation vector $u_{i,t}=(u_{i,t,1},\ldots,\\ u_{i,t,d})^{\top}\in{\mathbb{R}^{d}}$ is chosen independently by each agent $i\in\mathcal{N}$ from others and previous samples. In addition, the elements of $u_{i,t}$ are assumed i.i.d. with $\mathbb{E}(u_{i,t,d})^{2}=\sigma_{1}>0$, which implies that $\mathbb{E}[u_{i,t}u_{i,t}^{T}]=\sigma_{1}I_{d}$. Then, we further assume that there exists a constant $\sigma_{2}>0$ such that $\|u_{i,t}\|\leq\sigma_{2}.$
Remark 1
An example of a perturbation vector satisfying Assumption 5, we can choose every dimension of $u_{i,t}$ from the symmetric Bernoulli distribution on $\{-\tfrac{1}{\sqrt{d}},\tfrac{1}{\sqrt{d}}\}$. Then, $\sigma_{1}=\frac{1}{d},\sigma_{2}=1$.
Lemma 1
(Proposition $3.$ of mhanna2023single) Under Assumptions 4 and 5, then, $g_{i,t}$ is a biased estimator of the agent’s gradient $\nabla{F_{i}}({x_{i,t}}),\forall i\in{\mathcal{V}}$ for every $t\geq 0$, i.e.,
$\mathbb{E}{{}_{u,\xi,\varphi}}[{g_{i,t}}|{{\mathcal{H}}_{t}}]=d{\sigma_{1}}[\nabla{F_{i}}({x_{i,t}})+{b_{i,t}}],$ (7)
with
${{b_{i,t}}}={\frac{{{\gamma_{g}}}}{{2{\sigma_{1}}}}\mathbb{E}{{}_{u}}[{u_{i,t}}u_{i,t}^{T}{\nabla^{2}}{F_{i}}(\upsilon){u_{i,t}}|{{\mathcal{H}}_{t}}}$], (8)
where ${{b_{i,t}}}\in\mathbb{R}^{d}$ is the bias with respect to the true gradient, and $\upsilon\in[x_{i,t},\,x_{i,t}+\gamma_{g}u_{i,t}]$. Moreover, ${{\mathcal{H}}_{t}}=\{{{\mathbf{x}_{0}}},{\xi_{0}},...,{{\mathbf{x}_{t-1}}},{\xi_{t-1}},{{\mathbf{x}_{t}}}\}$ denotes the history sequence, where $\mathbf{x}_{t}$}} is defined in Section 3.2.
2.4 Compression operators
To reduce communication overhead, we consider a scenario where information exchange between agents is compressed. Specifically, we consider the compressors $\mathcal{C}(\cdot)$ with bounded relative compression error satisfying the following assumption.
Definition 1 (Compression)
A function $\mathcal{C}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is called a compression operator, if there exists a positive constant $\omega\in(0,1]$, such that for every $x\in\mathbb{R}^{d}$:
$\mathbb{E}_{\mathcal{C}}[\|x-\mathcal{C}(x)\|_{2}^{2}]\leq(1-\omega)\|x\|_{2}^{2},$ (9)
where expectation is taken over the randomness of $\mathcal{C}$. We assume that $\mathcal{C}(0)=0$.
Remark 2
As stated in liao2022compressed is naturally satisfied by various compressors in distributed algorithms. For example, the Randk sparsifier (beznosikov2023biased), the Topk sparsifier (zou2022downlink) and norm-sign compressor (yi2022communication).
3 Main results
In this section, we propose a Compressed Momentum-based Single-point Zeroth-Order (CMSGD_\_1P) algorithm. First, we introduce the CMSGD_\_1P algorithm and its motivation in Section 3.1. Then, we provide the supporting lemmas and establish the convergence of the CMSGD_\_1P algorithm in Section 3.2.
3.1 Algorithm Description
When the gradient information is available, in yu2019linear, the following distributed first-order momentum-based algorithm was proposed to solve (1):
$m_{i,t}$ $=\beta m_{i,{t-1}}+\nabla F_{i}(x_{i,t}),$ (10a)
$x_{i,{t}}$ $=x_{i,{t-1}}-\eta m_{i,t},$ (10b)
where $m_{i,t}$ is the local momentum of agent $i$; $\eta$ means SGD step-size; $\beta$ denotes momentum coefficient.
More notably, the first-order information is available, which may not always hold in practice, as gradient computation can be impractical or too costly for large-scale or highly complex systems. Then, we replace the gradient in (10a) with the stochastic gradient estimator (6).
To realize exchange of compressed parameters between agents, each agent $i$ maintains an estimate $\hat{x}_{i,t}$ of $x_{i,t}$ at time step $t$. Specifically, the updates for agent $i\in\mathcal{V}$ can be formulated as follows:
${\hat{x}_{j,t+1}}={\hat{x}_{j,t}}+{\mathcal{C}}({x_{i,t+\frac{1}{2}}}-{\hat{x}_{i,t}}),$ (11)
with
${x_{i,t+\frac{1}{2}}}={x_{i,t}}-\eta(\beta{m_{i,t}}+{g_{i,t}}),$ (12)
where $x_{i,t+\frac{1}{2}}$} is formed by every agent computes a stochastic gradient $g_{i,t}$ and takes a local SGD step on $x_{i,t}$ using momentum-based method.
Based on (6), (10a), (11) and (12), we summarize the algorithm updates in Algorithm 1.
Algorithm 1 CMSGD_1P Algorithm
1:Input: Stopping time $T$, adjacency matrix $\mathbf{W}$, and positive parameters $\gamma_{x}$, $\eta$, $\beta$, $\gamma_{g}$, $u_{i,t}$.
2:Initialization: For all $i\in\mathcal{V}$, initial variable $x_{i,0}\in\mathbb{R}^{d}$, ${x}_{i,-1}:=\textbf{0}$, $\hat{x}_{i,-1}:=\textbf{0}$, $m_{i,-1}=\textbf{0}\in\mathbb{R}^{d}$.
3:for $t=0$ to T1T-1 in parallel on all agents $i$ do
4:  One-point zeroth-order gradient:
  • Sample $\xi_{i,t}$ from the distribution $\mathcal{D}_{i}$;
  • Sample $u_{i,t}$ from the symmetrical distribution;
  • Query $f_{i}(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t})$;
  • Compute stochastic gradient $g_{i,t}$ using (6).
5:  Update $m_{i,t}$ by
${m_{i,t}}=\beta{m_{i,t-1}}+{g_{i,t}}.$ (13)
6:  Update $x_{i,t+\frac{1}{2}}$ according to (12).
7:  Compute $C(x_{i,t+\frac{1}{2}}-\hat{x}_{i,t})$ and broadcast it to its neighbors $\mathcal{N}_{i}$ according to (11).
8:  Consensus:
${x_{i,t+1}}={x_{i,t+\frac{1}{2}}}+{\gamma_{x}}\sum\limits_{j\in{{\mathcal{V}}_{i}}}{{w_{ij}}}({\hat{x}_{j,t+1}}-{\hat{x}_{i,t+1}}).$ (14)
9:end for
10:Output: $\{x_{i}(T)\}$.
3.2 Supporting Lemmas
For ease of presentation, we introduce the following notations:
${{\mathbf{x}_{t}}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{x_{1,t}^{T}},.,{x_{n,t}^{T}}]^{T}\in{\mathbb{R}^{nd}},{{\bar{\mathbf{x}}}_{t}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{{\bar{x}}_{t}^{T}},.,{{\bar{x}}_{t}^{T}}]^{T}\in{\mathbb{R}^{nd}},$
${{\hat{\mathbf{x}}_{t}}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{{\hat{x}}_{1,t}^{T}},.,{{\hat{x}}_{n,t}^{T}}]^{T}\in{\mathbb{R}^{nd}},{\mathbf{m}_{t}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{m_{1,t}^{T}},.,{m_{n,t}^{T}}]^{T}\in{\mathbb{R}^{nd}},$
${\mathbf{g}_{t}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{g_{1,t}^{T}},.,{g_{n,t}^{T}}]^{T}\in{\mathbb{R}^{nd}},$
$\mathcal{C}(\mathbf{x})$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[\mathcal{C}(x_{1})^{T},.,\mathcal{C}(x_{n})^{T}]^{T}\in{\mathbb{R}^{nd}},$
where ${\bar{x}}_{t}=\frac{1}{n}\sum\limits_{i=1}^{n}{{x_{i,t}}}$ denotes the average of node parameters.
To begin with, we consider the following Lyapunov candidate function.
${\chi_{t}}=\mathbb{E}\|{{\mathbf{x}_{t}}-{{\bar{\mathbf{x}}}_{t}}}\|_{F}^{2}+\mathbb{E}\|{{\mathbf{x}_{t}}-{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2},$ (15)
where incorporates the following error terms: the consensus error term $\mathbb{E}\|{{\mathbf{x}_{t}}-{{\bar{\mathbf{x}}}_{t}}}\|_{F}^{2}$ and the compressed error term $\mathbb{E}\|{{\mathbf{x}_{t}}-{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2}$. They play the key role for obtaining the convergence results of Algorithm 1.
Lemma 2
(Consensus error) Let $\mathbf{x}_{t}$} be the sequence generated by Algorithm 1, we have the following bound
$\mathbb{E}$ $\|\mathbf{x}_{t}-\bar{\mathbf{x}}_{t}\|_{F}^{2}$ (16)
$\leq$ ${\varepsilon_{1}}(1+\alpha_{2}^{-1})\mathbb{E}\|{{{\bar{\mathbf{x}}}_{t-1}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}+{{{{\varepsilon_{4}}\mathbb{E}}}{\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}}}$
$+[(1+{\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})+{\varepsilon_{3}}]\mathbb{E}\|{\eta(\beta{{\mathbf{m}}_{t-1}}+{{\mathbf{g}}_{t-1}})}\|_{F}^{2}.$
where $\alpha_{2}>0$ and $\varepsilon_{1}$$\varepsilon_{4}$} are given in Appendix 6.2.
Proof. See Appendix 6.2 for the proof.
Lemma 3
(Compressed error) Let $\mathbf{x}_{t}$} be the sequence generated by Algorithm 1, we have the following bound
$\mathbb{E}\|{{{\mathbf{x}}_{t}}-{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2}\leq{\kappa_{3}}\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\bar{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$ (17)
$+\{(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}})+{\kappa_{2}}\}\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$
$+\{(1+{\alpha_{5}})(1-\omega)(1+\alpha_{8}^{-1})+{\kappa_{1}}\}\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}.$
where $\alpha_{5},\alpha_{8}>0$ and $\kappa_{1}$$\kappa_{3}$} are provided in Appendix 6.2.
Proof. See Appendix 6.2 for the proof.
Lemma 4
Suppose Assumptions 45 hold, one has
$\mathbb{E}$ $\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}\leq\frac{{16{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(1-\beta)\gamma_{g}^{2}}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}$ (18)
$+\frac{{16{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}{\chi_{t-1}}+2{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}),$
where $\kappa_{4}$ can be found in the proof of Lemma 4.
Proof. See Appendix 6.2 for the proof.
Now, we show the convergence of CMSGD_\_1P algorithm under general nonconvex settings.
Theorem 1
Suppose Assumptions 15 hold. For the total number of iterations $T\geq\max\{{c_{1}},{c_{2}},{c_{3}},{c_{4}},{c_{5}},{c_{6}},{c_{7}}\}$, each agent runs Algorithm 1 with
${\gamma_{x}}\in(max\{{m_{1}},{m_{3}}\},min\{{m_{2}},{m_{4}}\}),$
$\gamma_{g}\geq\max\{{{\tilde{\gamma}}_{1}},{{\tilde{\gamma}}_{2}},{{\tilde{\gamma}}_{3}}\},\penalty 10000\ \eta=(1-\beta)\sqrt{\frac{n}{T}},$
where $m_{1}$$m_{4}$} and $c_{1}$$c_{7}$}, $\rho,\rho_{0},\rho_{2}$, $\rho_{3}$ and ${\tilde{\gamma}}_{1}$${\tilde{\gamma}}_{3}$} are provided in Appendix 6.3. Then,
$\frac{1}{T}$ $\sum\limits_{t=0}^{T-1}{\sum\limits_{i=1}^{n}{\mathbb{E}{{\|{{x_{i,t}}-{{\bar{x}}_{t}}}\|}^{2}}}}={\mathcal{O}}(\frac{1}{{T}})+{\mathcal{O}}(\frac{1}{{T}\gamma_{g}^{2}}),$ (19)
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}{{\|{\nabla{f}({{\bar{x}}_{t}})}\|}^{2}}$ $={\mathcal{O}}(\frac{1}{{\sqrt{nT}}})+{\mathcal{O}}(\frac{{\sqrt{n}}}{{\sqrt{T}}})+{\mathcal{O}}(\frac{{\sqrt{n}}}{{\sqrt{T}\gamma_{g}^{2}}})$
$+{\mathcal{O}}(\frac{{\sqrt{{n^{3}}}}}{{\sqrt{{T^{3}}}\gamma_{g}^{2}}})+{\mathcal{O}}(\frac{{\sqrt{{n^{3}}}}}{{\sqrt{{T^{3}}}\gamma_{g}^{4}}})+{\mathcal{O}}({{\gamma_{g}^{2}}}).$ (20)
Proof: The detailed proof is given in Appendix 6.3.
Remark 3
It is worth noting that the inherent noise of the single-point zero-order estimator and the use of constant step sizes, the RHS of (1) contains ${\mathcal{O}}({{\gamma_{g}^{2}}})$, which does not indicate a failure of Algorithm 1. Rather, it reflects the fact that under a noisy single-point estimator with constant step sizes, the algorithm can only converge to a neighborhood of a stationary point. The radius of this neighborhood is proportional to $\gamma_{g}$, $\sigma_{1}$, $\sigma_{2}$ and $\vartheta_{1}$.
From the right-hand side of (1), it follows immediately that sublinear convergence rate can be fulfilled if ${\gamma_{g}}={T^{-\frac{1}{8}}}$, which is presented in the following result.
Corollary 1
Under the same assumptions and parameters settings in Theorem 1, let ${\gamma_{g}}={T^{-\frac{1}{8}}}$ and $T\geq\max\{{c_{1}},{c_{2}},{\tilde{c}_{3}},{\tilde{c}_{4}},{\tilde{c}_{5}},{\tilde{c}_{6}},{\tilde{c}_{7}}\}$, then,
$\frac{1}{T}$ $\sum\limits_{t=0}^{T-1}{\sum\limits_{i=1}^{n}{\mathbb{E}{{\|{{x_{i,t}}-{{\bar{x}}_{t}}}\|}^{2}}}}={\mathcal{O}}(\frac{1}{{T}})+{\mathcal{O}}(\frac{1}{{\sqrt[4]{{{T^{3}}}}}}),$ (21)
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}{{\|{\nabla{f}({{\bar{x}}_{t}})}\|}^{2}}=$ ${\mathcal{O}}(\frac{1}{{\sqrt{nT}}})+{\mathcal{O}}(\frac{{\sqrt{n}}}{{\sqrt{T}}})+{\mathcal{O}}(\frac{{\sqrt{n}}}{{\sqrt[4]{T}}})$
$+{\mathcal{O}}(\frac{{\sqrt{{n^{3}}}}}{{\sqrt[4]{{{T^{5}}}}}})+{\mathcal{O}}(\frac{{\sqrt{{n^{3}}}}}{T}),$ (22)
where
${{\tilde{c}}_{3}}$ $={(\frac{{320n{{(1-\beta)}^{2}}{\varepsilon_{0}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(16+5\omega)}})^{\frac{4}{3}}},$
${\tilde{c}_{4}}$ $={(\frac{{720n{\lambda^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}{{(1-\beta)}^{2}}}}{{\omega{\delta^{2}}}})^{\frac{4}{3}}},$
${\tilde{c}_{5}}$ $={(40n{(1-\beta)^{2}}{\varepsilon_{0}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2})^{\frac{4}{3}}},$
${\tilde{c}_{6}}$ $={(\frac{{192n{{(1-\beta)}^{2}}{\omega^{3}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\varepsilon_{5}^{3}{\varepsilon_{6}}}})^{\frac{4}{3}}},$
${\tilde{c}_{7}}$ $={(\frac{{16n{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}{{(1-\beta)}^{2}}}}{{{{(\frac{{\delta{\varepsilon_{8}}}}{{30}})}^{\frac{3}{2}}}}})^{\frac{4}{3}}}.$
4 Simulations
In this section, we provide the results of the simulations to prove the effectiveness of our algorithm.
First, the system consists of $n=6$ agents and dimension $d=30$ that exchange information through a connected undirected communication graph. And our theoretical results via numerical experiments on a nonconvex distributed binary classification task (xie2024communication):
$f_{i}(x,\xi_{i})={\hbar_{1}}(x)+{\hbar_{2}}(x),$ (23)
where ${\hbar_{1}}(x)=\sum_{j=1}^{m_{i}}\log(1+e^{-p_{ij}x^{\top}q_{ij}})$ denotes the logistic regression function, $m_{i}=200$ denotes the number of local observations per agent. $\xi_{i}=(p_{ij},q_{ij})$, $p_{ij}\in\mathbb{R}^{d}$ stands for the Gaussian features regenerated per iteration, $q_{ij}\in\{-1,1\}$ is labels. ${\hbar_{2}}(x)=\sum_{\iota=1}^{n}\frac{\varpi\kappa[x]_{\iota}^{2}}{1+\kappa[x]_{\iota}^{2}}$ represents a nonconvex regularizer with the regularization parameters $\varpi=0.001$ and $\kappa=1$. $x]_{\iota}$ is the $\iota$-th coordinate of $x\in\mathbb{R}^{d}$.
Moreover, we adopt the unbiased 2-bit quantizer in wang2025compressed, which satisfies Definition 1. The quantization operator $\mathcal{C}(\cdot)$ takes the form
$\mathcal{C}(x)=\frac{\|x\|_{\infty}}{2^{{k_{2}}-1}}\mathrm{sgn}(x)\circ\lfloor\frac{2^{{{k_{2}}}-1}|x|}{\|x\|_{\infty}}+{\varpi_{1}}\rfloor,$ (24)
where ${k_{2}}=2$, $\varpi_{1}$ is the random perturbation vector uniformly chosen from $0,1]^{d}$.
The performance metric
$P(T)=\min_{t\in T}\{\mathbb{E}_{\xi}\|\nabla f(\bar{x}_{t},\xi)\|^{2}+\tfrac{1}{n}\sum_{i=1}^{n}\|x_{i,t}-\bar{x}_{t}\|^{2}\}$
is evaluated in terms of both the iteration rounds and the communication bits among agents. The hyperparameters used in the experiments are summarized in Table 1.
We compare the Algorithm 1 against distributed stochastic gradient tracking method with a one-point gradient estimator (DSGT_\_1P) in mhanna2023single and one-point zeroth-order stochastic compressed distributed primal-dual algorithm (DZOS_\_PDA).
Fig. 1 shows that CMSGD_\_1P converges slightly slower than DSGT_\_1P but still faster than DZOS_\_PDA. Furthermore, Fig. 2 illustrates that compared to other algorithms, CMSGD_\_1P converges to the same accuracy with fewer bits. This also demonstrates the efficiency of CMSGD_\_1P.
Table 1: Hyperparameters for Algorithm 1
Algorithm $\gamma_{x}$ $\eta$ $\beta$ $\gamma_{g}$
CMSGD_\_1P $0.3$ $0.02$ $0.9$ $0.99$
DZOS_\_PDA $0.3$ 0.02 - $0.99$
DSGT_\_1P - $0.02$ - $0.99$
Figure 1: Evolutions of $P(T)$ with respect to the number of iterations.
Figure 2: Evolutions of $P(T)$ with respect to the number of transmitted bits.
5 Conclusion
In this paper, we proposed the CMSGD_\_1P algorithm for stochastic distributed nonconvex optimization under the assumption that the gradient is not available and that only noisy single queries of the objective function are available at a time. We confirmed that the proposed algorithm achieves the convergence rate $\mathcal{O}(\frac{1}{\sqrt[4]{T}})$ with the exact solution under diminishing step sizes and can achieve the sublinear convergence rate ${\mathcal{O}}(\frac{1}{{\sqrt{T}}})$ towards a neighborhood of the stationary point under fixed step sizes. The simulation results imply that CMSGD_\_1P algorithm can achieve better performance than existing methods that do not incorporate momentum updates. Future directions include extending the algorithm to the one-point residual-feedback oracle or incorporating the Byzantine-robust aggregation rules.
6 APPENDIX
6.1 Useful Lemmas
The following results are used in the proofs. More notably, unless specified otherwise, for a vector $\mathbf{u}$, we write $\|\mathbf{u}\|$ to denote the $\ell_{2}$-norm $\|\mathbf{u}\|_{2}$.
Lemma 5 (singh2021squarm)
Consider any two matrices $\mathbf{A}\in\mathbb{R}^{m\times n}$, $\mathbf{B}\in\mathbb{R}^{n\times n}$. Then the following holds
$\|\mathbf{AB}\|_{F}\leq\|\mathbf{A}\|_{F}\|\mathbf{B}\|_{2}.$ (25)
Lemma 6 (koloskova2019decentralized)
For doubly stochastic matrix $\mathbf{W}$ with second largest eigenvalue $1-\delta=|\lambda_{2}(\mathbf{W})|<1$, we have
$\|\mathbf{W}-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}\|=(1-\delta).$ (26)
6.2 Proof of Lemmas 24
We denote the following notations:
${{\varepsilon_{1}}}=$ ${(1+{\alpha_{1}}){{(1-{\gamma_{x}}\delta)}^{2}}},$
${{\varepsilon_{2}}}=$ ${(1+\alpha_{1}^{-1})\gamma_{x}^{2}{\lambda^{2}}},$
$\varepsilon_{3}=$ $\varepsilon_{2}(1+\alpha_{2}^{-1})[(1+\alpha_{3})(1-\omega)(1+\alpha_{4}^{-1})+(1+\alpha_{3}^{-1})],$
$\varepsilon_{4}=$ ${\varepsilon_{2}}(1+\alpha_{2}^{-1})(1+{\alpha_{3}})(1-\omega)(1+{\alpha_{4}}),$
${\kappa_{1}}=$ $(1+\alpha_{5}^{-1})\gamma_{x}^{2}{\lambda^{2}}\{(1+{\alpha_{6}})(1+{\alpha_{7}})[(1+{\alpha_{3}})(1-\omega)$
$(1+\alpha_{4}^{-1})+(1+\alpha_{3}^{-1})]+(1+\alpha_{6}^{-1})\},$
${\kappa_{2}}=$ $(1+\alpha_{5}^{-1})\gamma_{x}^{2}{\lambda^{2}}(1+{\alpha_{6}})(1+{\alpha_{7}})(1+{\alpha_{3}})$
$(1-\omega)(1+{\alpha_{4}}),$
${\kappa_{3}}=$ $(1+\alpha_{5}^{-1})\gamma_{x}^{2}{\lambda^{2}}(1+{\alpha_{6}})(1+\alpha_{7}^{-1}),$
where the parameters $\alpha_{1}$$\alpha_{8}$ are positive constants.
To simplify the analysis, we present the compact form of (13)–(14) as follows.
${\mathbf{m}_{t}}=\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t}};$ (27a)
${\mathbf{x}_{t+\frac{1}{2}}}={\mathbf{x}_{t}}-\eta(\beta{\mathbf{m}_{t}}+{\mathbf{g}_{t}});$ (28a)
${\mathbf{c}_{t}}={\mathcal{C}}({\mathbf{x}_{t+\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t}});$ (28b)
${{\hat{\mathbf{x}}}_{t+1}}={{\hat{\mathbf{x}}}_{t}}+{\mathbf{c}_{t}};$ (28c)
${\mathbf{x}_{t+1}}={\mathbf{x}_{t+\frac{1}{2}}}+{\gamma_{x}}{{\hat{\mathbf{x}}}_{t+1}}(\mathbf{W}-\mathbf{I}).$ (28d)
Now it is ready to prove Lemmas 24.
A. The proof of Lemma 2
By virtue of the (28d) and the Assumption 3, we can conclude that $\bar{\mathbf{x}}_{t}={\bar{\mathbf{x}}}_{t-\tfrac{1}{2}}$. Thus,
$\|\mathbf{x}_{t}-\bar{\mathbf{x}}_{t}\|_{F}^{2}=$ $\|{{\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t}}+{\gamma_{x}}{{\hat{\mathbf{x}}}}_{t}}(\mathbf{W}-\mathbf{I})\|_{F}^{2}$
$=$ $\|(\mathbf{x}_{t-\frac{1}{2}}-\bar{\mathbf{x}}_{t-\frac{1}{2}})[(1-\gamma_{x})\mathbf{I}+\gamma_{x}\mathbf{W}]$
$+\gamma_{x}(\hat{\mathbf{x}}_{t}-\mathbf{x}_{t-\frac{1}{2}})(\mathbf{W}-\mathbf{I})\|_{F}^{2}$
$\mathop{\leq}\limits^{(a)}$ $(1+{\alpha_{1}})\|{({\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{\mathbf{x}}}_{t-\frac{1}{2}}})[(1-{\gamma_{x}})\mathbf{I}+{\gamma_{x}}\mathbf{W}]}\|_{F}^{2}$
$+(1+\alpha_{1}^{-1})\|\gamma_{x}(\hat{\mathbf{x}}_{t}-\mathbf{x}_{t-\frac{1}{2}})(\mathbf{W}-\mathbf{I})\|_{F}^{2}$
$\mathop{\leq}\limits^{(b)}$ $(1+{\alpha_{1}})\|{({\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{\mathbf{x}}}_{t-\frac{1}{2}}})[(1-{\gamma_{x}})\mathbf{I}+{\gamma_{x}}\mathbf{W}]}\|_{F}^{2}$
$+(1+\alpha_{1}^{-1})\gamma_{x}^{2}\|{{{\hat{\mathbf{x}}}_{t}}-{{{\mathbf{x}}}_{t-\frac{1}{2}}}}\|_{F}^{2}\|{\mathbf{W}-\mathbf{I}}\|_{2}^{2}$
$\mathop{\leq}\limits^{(c)}$ $(1+{\alpha_{1}})\|{({\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{\mathbf{x}}}_{t-\frac{1}{2}}})[(1-{\gamma_{x}})\mathbf{I}+{\gamma_{x}}\mathbf{W}]}\|_{F}^{2}$
$+(1+\alpha_{1}^{-1})\gamma_{x}^{2}\lambda^{2}\|{{{\hat{\mathbf{x}}}_{t}}-{{{\mathbf{x}}}_{t-\frac{1}{2}}}}\|_{F}^{2},$ (29)
where $(a)$, $(b)$ and $(c)$ hold due to Lemma 5 and $\lambda=\max_{i}\{1-\lambda_{i}(\mathbf{W})\}\Rightarrow\|\mathbf{W}-\mathbf{I}\|_{2}^{2}\leq\lambda^{2}$.
(i) With respect to $\|{({\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{\mathbf{x}}}_{t-\frac{1}{2}}})[(1-{\gamma_{x}})\mathbf{I}+{\gamma_{x}}\mathbf{W}]}\|_{F}$, we have
$\|{({\mathbf{x}_{t-\frac{1}{2}}}}$ $-{{{\bar{\mathbf{x}}}_{t-\frac{1}{2}}})}{[(1-{\gamma_{x}})\mathbf{I}+{\gamma_{x}}\mathbf{W}]}\|_{F}$
$\leq$ $(1-{\gamma_{x}}){\|{{\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}}}\|_{F}}+{\gamma_{x}}{\|{({\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}}){\bf{W}}}\|_{F}}$
$=$ $(1-{\gamma_{x}}){\|{{{{\mathbf{x}}}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}}}\|_{F}}$
$+{\gamma_{x}}{\|{({{{\mathbf{x}}}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}})({\bf{W}}-\frac{{{{\mathbf{11}}^{T}}}}{n})}\|_{F}}$
$\leq$ $(1-{\gamma_{x}}){\|{{{{\mathbf{x}}}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}}}\|_{F}}+{\gamma_{x}}(1-\delta){\|{{{{\mathbf{x}}}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}}}\|_{F}}$
$=$ $(1-{\gamma_{x}}\delta){\|{{{{\mathbf{x}}}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}}}\|_{F}},$ (30)
where the first equality holds due to $({{{{\mathbf{x}}}}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}})\frac{{{{\mathbf{11}}^{T}}}}{n}=\mathbf{0}$ and the last inequality holds based on Lemma 6.
(ii) In what follows, we show the upper bound of $\|{{\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}}}\|_{F}^{2}$. Taking the average of both sides of (28a), we have
${\mathbf{\bar{x}}_{t-\frac{1}{2}}}={\mathbf{\bar{x}}_{t-1}}-\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})\frac{{{{\mathbf{11}}^{T}}}}{n},$ (31)
From (6.2), we have
$\mathbb{E}\|{\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{{\mathbf{x}}}}_{t-\frac{1}{2}}}\|_{F}^{2}$ (32)
$\leq\mathbb{E}\|{{\bar{\mathbf{x}}}_{t-1}}-{{{{\mathbf{x}}}}_{t-1}}-\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})(\frac{{{{\mathbf{11}}^{T}}}}{n}-\mathbf{I})\|_{F}^{2}.$
Similarly, from (28a) we have
$\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{{{\mathbf{x}}}_{t-\frac{1}{2}}}}\|_{F}^{2}\leq\mathbb{E}\|{{{\hat{\mathbf{{x}}}}_{t}}-{\mathbf{x}_{t-1}}+\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}.$ (33)
By substituting (6.2), (31), (33) into (6.2) and taking the expectation of both sides yields that
$\mathbb{E}\|\mathbf{x}_{t}$ $-\bar{\mathbf{x}}_{t}\|_{F}^{2}$
$\leq$ ${{\varepsilon_{1}}}\mathbb{E}\|{{\mathbf{x}_{t-\frac{1}{2}}}-{{\bar{\mathbf{x}}}_{t-\frac{1}{2}}}}\|_{F}^{2}+{{\varepsilon_{2}}}\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{{\mathbf{x}}_{t-\frac{1}{2}}}}\|_{F}^{2}$
$\leq$ ${\varepsilon_{1}}\mathbb{E}\|{{{\bar{\mathbf{x}}}_{t-1}}-{{\mathbf{x}}_{t-1}}-\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})(\frac{{\mathbf{1}{\mathbf{1}^{T}}}}{n}-\mathbf{I})}\|_{F}^{2}$
$+{\varepsilon_{2}}\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{{\mathbf{x}}_{t-1}+\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}}\|_{F}^{2},$
$\leq$ ${\varepsilon_{1}}(1+\alpha_{2}^{-1})\mathbb{E}\|{{{\bar{\mathbf{x}}}_{t-1}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}$
$+{\varepsilon_{1}}(1+{\alpha_{2}})\mathbb{E}\|{\eta(\beta{{\mathbf{m}}_{t-1}}+{{\mathbf{g}}_{t-1}})(\frac{{{{\mathbf{11}}^{T}}}}{n}-I)}\|_{F}^{2}$
$+{\varepsilon_{2}}(1+\alpha_{2}^{-1})\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}$
$+{\varepsilon_{2}}(1+{\alpha_{2}})\mathbb{E}\|{\eta(\beta{{\mathbf{m}}_{t-1}}+{{\mathbf{g}}_{t-1}})}\|_{F}^{2}$
$\leq$ ${\varepsilon_{1}}(1+\alpha_{2}^{-1})\mathbb{E}\|{{{\bar{\mathbf{x}}}_{t-1}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}$
$+{\varepsilon_{2}}(1+\alpha_{2}^{-1})\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}$
$+(1+{\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})\mathbb{E}\|{\eta(\beta{{\mathbf{m}}_{t-1}}+{{\mathbf{g}}_{t-1}})}\|_{F}^{2},$ (34)
where the last equality holds due to ${\|{\frac{{\mathbf{1}{\mathbf{1}^{T}}}}{n}-\mathbf{I}}\|_{2}}=1$.
An important observation is that
$\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}}$ $-{{\mathbf{x}_{t-1}}}\|_{F}^{2}$ (35)
$\mathop{=}\limits^{(a)}$ $\mathbb{E}{\|{{\mathbf{x}}_{t-1}}-({{\hat{\mathbf{x}}}_{t-1}}+{\mathcal{C}}({{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}}))\|_{F}^{2}}$
$=$ $\mathbb{E}{\|{{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}}-{\mathcal{C}}({{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}})-{{\mathbf{x}}_{t-\frac{1}{2}}}+{{\mathbf{x}}_{t-1}}}\|_{F}^{2}}$
$\mathop{\leq}\limits^{(b)}$ ${\rm{(1+}}{\alpha_{3}}{\rm{)(1}}-\omega{\rm{)\mathbb{E}}}{\|{{\mathbf{x}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}}$
$+(1+\alpha_{3}^{-1})\mathbb{E}{\|{{\mathbf{x}_{t-1}}-{\mathbf{x}_{t-\frac{1}{2}}}}\|_{F}^{2}}$
$=$ ${\rm{(1+}}{\alpha_{3}}{\rm{)(1}}-\omega{\rm{)\mathbb{E}}}\|{{\mathbf{x}_{t-\frac{1}{2}}}-{{\mathbf{x}}}_{t-1}}+{{\mathbf{x}}}_{t-1}+{{\hat{\mathbf{x}}}_{t-1}}\|_{F}^{2}$
$+(1+\alpha_{3}^{-1})\mathbb{E}{\|{{\mathbf{x}_{t-1}}-{\mathbf{x}_{t-\frac{1}{2}}}}\|_{F}^{2}}$
$\mathop{\leq}\limits^{(c)}$ ${\rm{(1+}}{\alpha_{3}}{\rm{)(1}}-\omega{\rm{)[}}(1+\alpha_{4}^{-1}){\rm{\mathbb{E}}}{\|{{\mathbf{x}_{t-\frac{1}{2}}}-{\mathbf{x}_{t-1}}}\|_{F}^{2}}$
$+{\rm{(1+}}{\alpha_{4}}{\rm{)\mathbb{E}}}{\|{{\mathbf{x}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}}]$
$+(1+\alpha_{3}^{-1})\mathbb{E}{\|{{\mathbf{x}_{t-1}}-{{\mathbf{x}}_{t-\frac{1}{2}}}}\|_{F}^{2}}$
$=$ $[(1+{\alpha_{3}})(1-\omega)(1+\alpha_{4}^{-1})+(1+\alpha_{3}^{-1}){\rm{]\mathbb{E}}}{\|{{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}}$
$+(1+{\alpha_{3}})(1-\omega){\rm{(1+}}{\alpha_{4}}{\rm{)\mathbb{E}}}{\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}},$
where $(a)$ holds due to (28b)–(28c); $(b)$ holds the Definition 1; $(c)$ holds Lemma 5.
Consequently, substituting (35) into (6.2), one obtains
$\mathbb{E}\|\mathbf{x}_{t}$ $-\bar{\mathbf{x}}_{t}\|_{F}^{2}\leq{\varepsilon_{1}}(1+\alpha_{2}^{-1})\mathbb{E}\|{{{\bar{\mathbf{x}}}_{t-1}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}$
$+{\varepsilon_{3}}{{\mathbb{E}}{\|{{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}}}+{{{{\varepsilon_{4}}\mathbb{E}}}{\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}}}$
$+(1+{\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})\mathbb{E}\|{\eta(\beta{{\mathbf{m}}_{t-1}}+{{\mathbf{g}}_{t-1}})}\|_{F}^{2}.$ (36)
By virtue of (28a), it follows that
$\mathbb{E}{\|{{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}}=\mathbb{E}{\|\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})\|_{F}^{2}}.$ (37)
From (37), (6.2) can be rearranged as (16).
B. The proof of Lemma 3
From (28c)–(28d), we have
$\mathbb{E}\|{{{\mathbf{x}}_{t}}}$ $-{{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2}=\mathbb{E}\|{{{\mathbf{x}}_{t}}-({{\hat{\mathbf{x}}}_{t-1}}+{\mathcal{C}}({{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}}))}\|_{F}^{2}$
$=$ $\mathbb{E}\|{{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}}-{\mathcal{C}}({{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}})+{\mathbf{x}}_{t}}-{{\mathbf{x}}_{t-\frac{1}{2}}}\|_{F}^{2}$
$\leq$ $(1+{\alpha_{5}})(1-\omega)\mathbb{E}\|{{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$
$+(1+\alpha_{5}^{-1})\mathbb{E}\|{{{\mathbf{x}}_{t}}-{{\mathbf{x}}_{t-\frac{1}{2}}}}\|_{F}^{2}.$ (38)
(i) Now, we provide the upper bound of $\mathbb{E}\|{{{\mathbf{x}}_{t}}-{{\mathbf{x}}_{t-\frac{1}{2}}}}\|_{F}^{2}$. Based on the (28d), one has
$\mathbb{E}\|{{{\mathbf{x}}_{t}}}$ ${-{{\mathbf{x}}_{t-\frac{1}{2}}}}\|_{F}^{2}=\gamma_{x}^{2}\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}(\mathbf{W}-\mathbf{I})}\|_{F}^{2}$
$\mathop{=}\limits^{(a)}$ $\gamma_{x}^{2}\mathbb{E}\|{({{\hat{\mathbf{x}}}_{t}}-{{\bar{\mathbf{x}}}_{t-\frac{1}{2}}})(\mathbf{W}-\mathbf{I})}\|_{F}^{2}$
$\mathop{\leq}\limits^{(b)}$ $\gamma_{x}^{2}{\lambda^{2}}\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{{\bar{\mathbf{x}}}_{t-\frac{1}{2}}}}\|_{F}^{2}$
$=$ $\gamma_{x}^{2}{\lambda^{2}}\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{\mathbf{\bar{x}}_{t-1}}+\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})\frac{{{{\mathbf{11}}^{T}}}}{n}}\|_{F}^{2}$
$\leq$ $\gamma_{x}^{2}{\lambda^{2}}(1+{\alpha_{6}})\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{\mathbf{\bar{x}}_{t-1}}}\|_{F}^{2}+\gamma_{x}^{2}{\lambda^{2}}(1+\alpha_{6}^{-1})$
$\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})\frac{{{{\mathbf{11}}^{T}}}}{n}}\|_{F}^{2}$
$=$ $\gamma_{x}^{2}{\lambda^{2}}(1+{\alpha_{6}})\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{\mathbf{\bar{x}}_{t-1}}}\|_{F}^{2}+\gamma_{x}^{2}{\lambda^{2}}(1+\alpha_{6}^{-1})$
$\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}}})\|_{F}^{2}$
$\mathop{\leq}\limits^{(c)}$ $\gamma_{x}^{2}{\lambda^{2}}(1+{\alpha_{6}})[{(1+{\alpha_{7}})\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{{{\mathbf{x}}}_{t-1}}}\|_{F}^{2}}$
$+(1+\alpha_{7}^{-1})\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\bar{\mathbf{x}}}_{t-1}}}\|_{F}^{2}]$
$+\gamma_{x}^{2}{\lambda^{2}}(1+\alpha_{6}^{-1})\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}$
$=$ $\gamma_{x}^{2}{\lambda^{2}}{(1+{\alpha_{6}})(1+{\alpha_{7}})\mathbb{E}\|{{{\hat{\mathbf{x}}}_{t}}-{{{\mathbf{x}}}_{t-1}}}\|_{F}^{2}}$
$+\gamma_{x}^{2}{\lambda^{2}}(1+{\alpha_{6}})(1+\alpha_{7}^{-1})\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\bar{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$
$+\gamma_{x}^{2}{\lambda^{2}}(1+\alpha_{6}^{-1})\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2},$ (39)
where $(a)$ holds due to ${{\bar{\mathbf{x}}}_{t-\frac{1}{2}}}(\mathbf{W}-\mathbf{I})=\mathbf{0}$. The inequality $(b)$ follows $\lambda=\max_{i}\{1-\lambda_{i}(\mathbf{W})\}\Rightarrow\|\mathbf{W}-\mathbf{I}\|_{2}^{2}\leq\lambda^{2}$ and the validity of $(c)$ follows from Lemma 5.
By combining (6.2) with (32), one has
$\mathbb{E}$ $\|{{{\mathbf{x}}_{t}}-{{\mathbf{x}}_{t-\frac{1}{2}}}}\|_{F}^{2}$
$\leq$ ${\gamma_{x}^{2}}{\lambda^{2}}{(1+{\alpha_{6}})(1+{\alpha_{7}})[(1+{\alpha_{3}})(1-\omega)(1+\alpha_{4}^{-1})}$
$+{(1+\alpha_{3}^{-1}){\rm{]\mathbb{E}}}{\|{{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}}}$
$+{\gamma_{x}^{2}{\lambda^{2}}{(1+{\alpha_{6}})(1+{\alpha_{7}})}}(1+{\alpha_{3}})(1-\omega)$
${\rm{(1+}}{\alpha_{4}}{\rm{)\mathbb{E}}}{\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}}$
$+\gamma_{x}^{2}{\lambda^{2}}(1+{\alpha_{6}})(1+\alpha_{7}^{-1})\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\bar{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$
$+\gamma_{x}^{2}{\lambda^{2}}(1+\alpha_{6}^{-1})\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}.$ (40)
(ii) Then, we provide the upper bound of $\mathbb{E}\|{{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$. Based on Lemma 5, it readily follows that
$\mathbb{E}\|$ ${{{\mathbf{x}}_{t-\frac{1}{2}}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}=\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-\eta(\beta\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})-{{\hat{\mathbf{x}}}_{t-1}}\|_{F}^{2}$ (41)
$\leq$ $(1+{\alpha_{8}})\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$
$+(1+\alpha_{8}^{-1})\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}}})\|_{F}^{2}.$
By combining (6.2) with (6.2), (41) can be rearranged as (17).
C. The proof of Lemma 4
To begin with, we denote following notations:
${\ell_{1}}=$ ${\varepsilon_{1}(1+\alpha_{2}^{-1})+\kappa_{3}},$
${\ell_{2}}=$ ${\varepsilon_{4}+(1+\alpha_{5})(1-\omega)(1+\alpha_{8})+\kappa_{2}},$
${{\ell_{3}}}=$ ${\varepsilon_{3}}+(1+{\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})+(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}})+{\kappa_{1}},$
${\kappa_{4}}=$ $\frac{{4n{d^{2}}\sigma_{2}^{2}}[L_{{f_{1}}}^{2}\gamma_{g}^{2}\sigma_{2}^{2}+\gamma_{1}^{2}+{\vartheta_{2}}]}{{\gamma_{g}^{2}}}.$
We define an auxiliary sequence $\tilde{x}_{i,t}$} for each agent $i$ as follows:
${\tilde{x}_{i,t}}={x_{i,t}}-\frac{{\eta{\beta^{\rm{2}}}}}{{1-\beta}}{m_{i,t-1}},$ (42)
Then, let ${\bar{x}}_{t}^{a}=\frac{1}{n}\sum\limits_{i=1}^{n}{{{\tilde{x}}_{i,t}}}$, (42) can be transformed as
${\bar{x}}_{t}^{a}={{\bar{x}}_{t}}-\frac{{\eta{\beta^{\rm{2}}}}}{{1-\beta}}\frac{1}{n}\sum\limits_{i=1}^{n}{{m_{i,t-1}}}.$ (43)
From (42)–(43), we have
${\bar{x}}_{t+1}^{a}$ $={{\bar{x}}_{t+1}}-\frac{{\eta{\beta^{\rm{2}}}}}{{1-\beta}}\frac{1}{n}\sum\limits_{i=1}^{n}{{m_{i,t}}}$
$={{\bar{x}}_{t}}-\frac{\eta}{n}\sum\limits_{i=1}^{n}{(\beta{m_{i,t}}+{g_{i,t}})}-\frac{{\eta{\beta^{\rm{2}}}}}{{1-\beta}}\frac{1}{n}\sum\limits_{i=1}^{n}{{m_{i,t}}}$
$={{\bar{x}}_{t}}-\frac{\eta}{n}\sum\limits_{i=1}^{n}{{g_{i,t}}}-\frac{1}{n}(\eta\beta+\frac{{\eta{\beta^{\rm{2}}}}}{{1-\beta}})\sum\limits_{i=1}^{n}{{m_{i,t}}}$
$={{\bar{x}}_{t}}-\frac{\eta}{n}\sum\limits_{i=1}^{n}{{g_{i,t}}}-\frac{1}{n}\frac{{\eta\beta}}{{1-\beta}}\sum\limits_{i=1}^{n}{(\beta{m_{i,t-1}}+{g_{i,t}})}$
$={{\bar{x}}_{t}}-(\eta+\frac{{\eta\beta}}{{1-\beta}})\frac{1}{n}\sum\limits_{i=1}^{n}{{g_{i,t}}}-\frac{{\eta{\beta^{2}}}}{{1-\beta}}\frac{1}{n}\sum\limits_{i=1}^{n}{{m_{i,t-1}}}$
$={\bar{x}}_{t}^{a}-\frac{\eta}{{1-\beta}}\frac{1}{n}\sum\limits_{i=1}^{n}{{g_{i,t}}},$ (44)
where the second, fourth equalities hold due to (13).
Then, we define ${\chi_{t}}=\mathbb{E}\|{{\mathbf{x}_{t}}-{{\bar{\mathbf{x}}}_{t}}}\|_{F}^{2}+\mathbb{E}\|{{\mathbf{x}_{t}}-{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2}$, one has
${\chi_{t}}=$ $\mathbb{E}\|{{\mathbf{x}_{t}}-{{\bar{\mathbf{x}}}_{t}}}\|_{F}^{2}+\mathbb{E}\|{{\mathbf{x}_{t}}-{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2}$ (45)
$\leq$ $[{\varepsilon_{1}}(1+\alpha_{2}^{-1})+{\kappa_{3}}]\mathbb{E}\|{{{\bar{\mathbf{x}}}_{t-1}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}$
$+[{\varepsilon_{3}}+(1+{\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})+(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}^{-1}})$
$+{\kappa_{1}}]\mathbb{E}\|{\eta(\beta{{\mathbf{m}}_{t-1}}+{{\mathbf{g}}_{t-1}})}\|_{F}^{2}$
$+[{\varepsilon_{4}}+(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}})+{\kappa_{2}}]\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$
$\leq$ $max\{{{\ell_{1}},{\ell_{2}}}\}{\chi_{t-1}}+{\ell_{3}}\mathbb{E}\|{\eta(\beta{{\mathbf{m}}_{t-1}}+{{\mathbf{g}}_{t-1}})}\|_{F}^{2}.$
By employing the bounded of the expected norm squared in (6), we have
$\mathbb{E}\|{{\mathbf{g}_{t}}}$ $\|_{F}^{2}=\sum\limits_{i=1}^{n}{\mathbb{E}{{\|{{g_{i,t}}}\|}^{2}}}$
$=$ $\sum\limits_{i=1}^{n}{\mathbb{E}{{\|{\frac{d}{{{\gamma_{g}}}}{u_{i,t}}[{{f_{i}}({x_{i,t}}+{\gamma_{g}}{u_{i,t}},{\xi_{i,t}})+{\varphi_{i,t}}}]\|^{2}}}}}$
$=$ $\frac{{{d^{2}}}}{{\gamma_{g}^{2}}}\sum\limits_{i=1}^{n}{\mathbb{E}\|{{u_{i,t}}[{f_{i}}({x_{i,t}}+{\gamma_{g}}{u_{i,t}},{\xi_{i,t}})-{f_{i}}({{\bar{x}}_{t}}+{\gamma_{g}}{u_{i,t}},{\xi_{i,t}})}}$
$+{f_{i}}({{\bar{x}}_{t}}+{\gamma_{g}}{u_{i,t}},{\xi_{i,t}})-{f_{i}}({{\bar{x}}_{t}},{\xi_{i,t}})+{f_{i}}({{\bar{x}}_{t}},{\xi_{i,t}})$
${{-{f_{i}}({x_{i,t}},{\xi_{i,t}})+{f_{i}}({x_{i,t}},{\xi_{i,t}})+{\varphi_{i,t}}]}\|^{2}}$
$\leq$ $\frac{{4{d^{2}}}}{{\gamma_{g}^{2}}}\sum\limits_{i=1}^{n}{\mathbb{E}[{{\|{{u_{i,t}}}\|}^{2}}(2L_{{f_{1}}}^{2}{\|x_{i,t}}-{{{\bar{x}}_{t}}\|{{}^{2}}+L_{{f_{1}}}^{2}\gamma_{g}^{2}\|{u_{i,t}}\|{{}^{2}}}}$
[…]
$\leq$ $\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}\sum\limits_{i=1}^{n}\mathbb{E}\|{{x_{i,t}}-{{\bar{x}}_{t}}\|{{}^{2}}}+\kappa_{4}$
$\leq$ $\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}\chi_{t}+\kappa_{4},$ (46)
where the first and second inequality holds due to Assumptions 1, 4 and 5.
Using ${{\bar{g}}_{t}}={\frac{1}{n}\sum\limits_{i=1}^{n}{{g_{i,t}}}}$, based on the Jensen’s inequality and the convexity of norms, one has
${\|{{{\bar{g}}_{t}}}\|^{2}}={\|{\frac{1}{n}\sum\limits_{i=1}^{n}{{g_{i,t}}}}\|^{2}}\leq\frac{1}{n}{\sum\limits_{i=1}^{n}{\|{{g_{i,t}}}\|}^{2}}.$ (47)
We can further conclude that
$\mathbb{E}{\|{{{\bar{g}}_{t}}}\|^{2}}\leq\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}\chi_{t}+\frac{{{\kappa_{4}}}}{n}.$ (48)
From the (27a) and $\mathbf{m}_{-1}=\mathbf{0}$, we can derive that
$\mathbf{m}_{t-1}=\sum_{k=0}^{t-1}\beta^{t-1-k}\mathbf{g}_{k}.$ (49)
From (49), we have
$\mathbb{E}\|{{{\bf{m}}_{t-1}}}\|_{F}^{2}$ $=\mathbb{E}\|{\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{{\bf{g}}_{k}}}\|_{F}^{2}$
$=\mathbb{E}\|{{\theta_{t-1}}\sum\limits_{k=0}^{t-1}{\frac{{{\beta^{t-1-k}}}}{{{\theta_{t-1}}}}}{{\bf{g}}_{k}}}\|_{F}^{2}$
$=\theta_{t-1}^{2}\mathbb{E}\|{\sum\limits_{k=0}^{t-1}{\frac{{{\beta^{t-1-k}}}}{{{\theta_{t-1}}}}}{{\bf{g}}_{k}}}\|_{F}^{2}$
$\leq{\theta_{t-1}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}\mathbb{E}\|{{{\bf{g}}_{k}}}\|_{F}^{2}$
$\leq\frac{1}{{1-\beta}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}\mathbb{E}\|{{{\bf{g}}_{k}}}\|_{F}^{2}$
$\leq\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(1-\beta)\gamma_{g}^{2}}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}\chi_{k}+\frac{{{\kappa_{4}}}}{{(1-\beta)^{2}}},$ (50)
where $\theta_{t-1}:=\sum_{k=0}^{t-1}\beta^{t-1-k}=\frac{1-\beta^{t}}{1-\beta}\leq\frac{1}{1-\beta}$ and ${\alpha_{k}}=\frac{{{\beta^{t-1-k}}}}{{{\theta_{t-1}}}}$, one has $\sum\limits_{k=0}^{t-1}{{\alpha_{k}}}=1$; the second inequality holds due to (6.2).
As is well known,
$\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}\leq 2{\eta^{2}}\mathbb{E}\|{\beta{\mathbf{m}_{t-1}}}\|_{F}^{2}+2{\eta^{2}}\mathbb{E}\|{{\mathbf{g}_{t-1}}}\|_{F}^{2}.$ (51)
By combining (6.2) with (6.2), (51) can be rearranged as (18).
6.3 Proof of Theorem 1
To simplify the analysis, we denote the following notations:
${\varepsilon_{0}}$ $={1-\frac{\omega}{4}+\frac{{9{\lambda^{2}}}}{{{\delta^{2}}}}},$
${\varepsilon_{1}}$ $=\delta-\frac{{45\rho{\lambda^{2}}}}{{\omega\delta}},$
${\varepsilon_{2}}$ $=\frac{4}{5}-\rho{\varepsilon_{0}},$
${\varepsilon_{4}}$ $=\frac{4}{5}+\frac{\omega}{4}-\rho{\varepsilon_{0}},$
${\varepsilon_{5}}$ $=\frac{{\frac{2}{5}+\frac{\omega}{4}}}{{2(\frac{{3{\lambda^{2}}}}{\delta}+\frac{{18{\lambda^{2}}}}{{\omega\delta{\varepsilon_{0}}}})}},$
${{\varepsilon_{6}}}$ $={\delta^{3}}{\omega^{3}}+128\delta\omega{\lambda^{2}},$
$\rho$ $=\frac{{16{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}},$
${\rho_{0}}$ $=\frac{2}{{5{\varepsilon_{0}}}},$
${\rho_{1}}$ $=\frac{{10}}{9}(max\{{\ell_{1}},{\ell_{2}}\}+{\ell_{3}}\rho),$
${\rho_{2}}$ $=\frac{{\varepsilon_{5}^{3}{{\varepsilon_{6}}}}}{{12{\omega^{3}}}},$
${\varepsilon_{7}}$ $={{\varepsilon_{6}}}+1500{\rho_{0}}\delta{\lambda^{2}}+25{\omega^{2}}{\rho_{0}}\delta{\lambda^{2}},$
${\varepsilon_{8}}$ $=\sqrt[3]{{\frac{{12{\omega^{3}}}}{{{\varepsilon_{7}}}}}},$
${\rho_{3}}$ $={(\frac{{\delta{\varepsilon_{8}}}}{{30}})^{\frac{3}{2}}},$
${\varepsilon_{9}}$ $={\varepsilon_{6}}+1500\rho\delta\lambda^{2}+25\omega^{2}\rho\delta\lambda^{2},$
$\kappa_{5}$ $=100\delta\omega\lambda^{2}+1500\rho\delta\lambda^{2}+25\delta\omega^{2}\rho\lambda^{2}.$
Now, we provide the upper bound of $\frac{1}{T}\sum\limits_{t=0}^{T-1}{{\chi_{t}}}$. From (45), (6.2),
(6.2) and (51), one has
$\chi_{t}\leq$ $max\{{\ell_{1}},{\ell_{2}}\}{\chi_{t-1}}+{\ell_{3}}(\frac{{16{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(1-\beta)\gamma_{g}^{2}}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}$ (52)
$+\frac{{16{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}{\chi_{t-1}}+2{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}))$
$=$ $(max\{{\ell_{1}},{\ell_{2}}\}+\frac{{16{\ell_{3}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}){\chi_{t-1}}$
$+\frac{{16\ell_{3}}{{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(1-\beta)\gamma_{g}^{2}}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}$
$+2{\ell_{3}}{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}).$
Taking summation from $t=0$ to T1T-1, we get
$\sum\limits_{t=0}^{T-1}$ ${{\chi_{t}}}\leq(max\{{\ell_{1}},{\ell_{2}}\}+\frac{{16{\ell_{3}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}})\sum\limits_{t=0}^{T-1}{{\chi_{t-1}}}$
$+\frac{{16{\ell_{3}}{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(1-\beta)\gamma_{g}^{2}}}\sum\limits_{t=0}^{T-1}{\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}}$
$+2T{\ell_{3}}{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}).$ (53)
More notably,
$\sum_{j=0}^{T-2}\sum_{k=0}^{j}\beta^{\,j-k}\chi_{k}=\sum_{k=0}^{T-2}\sum_{j=k}^{T-2}\beta^{\,j-k}\chi_{k}.$ (54)
We prove (54) by mathematical induction on $T$.
(i) For $T=2$, we have
$\sum_{j=0}^{2-2}\sum_{k=0}^{j}\beta^{\,j-k}\chi_{k}=\sum_{j=0}^{0}\sum_{k=0}^{j}\beta^{\,j-k}\chi_{k}=\sum_{j=0}^{0}\sum_{k=0}^{0}\beta^{\,0}\chi_{0}=\chi_{0},$
and
$\sum_{k=0}^{2-2}\sum_{j=k}^{2-2}\beta^{\,j-k}\chi_{k}=\sum_{k=0}^{0}\sum_{j=k}^{0}\beta^{\,j-k}\chi_{k}=\sum_{k=0}^{0}\sum_{j=0}^{0}\beta^{\,0}\chi_{0}=\chi_{0}.$
Then, (54) holds for $T=2$.
(ii) Assume that (54) holds for $T\geq 2$, i.e.,
$\sum_{j=0}^{T-2}\sum_{k=0}^{j}\beta^{\,j-k}\chi_{k}=\sum_{k=0}^{T-2}\sum_{j=k}^{T-2}\beta^{\,j-k}\chi_{k}.$ (55)
(iii) We show that (54) also holds for $T+1$, i.e.,
$\sum_{j=0}^{(T+1)-2}\sum_{k=0}^{j}\beta^{\,j-k}\chi_{k}=\sum_{k=0}^{(T+1)-2}\sum_{j=k}^{(T+1)-2}\beta^{\,j-k}\chi_{k},$ (56)
which is
$\sum_{j=0}^{T-1}\sum_{k=0}^{j}\beta^{\,j-k}\chi_{k}=\sum_{k=0}^{T-1}\sum_{j=k}^{T-1}\beta^{\,j-k}\chi_{k}.$
For the left-hand side of (56), we have
$\sum_{j=0}^{T-1}\sum_{k=0}^{j}\beta^{\,j-k}\chi_{k}$ $=\sum_{j=0}^{T-2}\sum_{k=0}^{j}\beta^{\,j-k}\chi_{k}+\sum_{k=0}^{T-1}\beta^{\,(T-1)-k}\chi_{k}.$
From (55), one obtains
$\sum_{j=0}^{T-1}\sum_{k=0}^{j}$ $\beta^{\,j-k}\chi_{k}=\sum_{k=0}^{T-2}\sum_{j=k}^{T-2}\beta^{\,j-k}\chi_{k}+\sum_{k=0}^{T-1}\beta^{\,(T-1)-k}\chi_{k}$
$=\sum_{k=0}^{T-2}(\sum_{j=k}^{T-2}\beta^{\,j-k}\chi_{k}+\beta^{\,(T-1)-k}\chi_{k})+\beta^{\,0}\chi_{T-1}$
$=\sum_{k=0}^{T-2}\sum_{j=k}^{T-1}\beta^{\,j-k}\chi_{k}+\chi_{T-1}$
$=\sum_{k=0}^{T-1}\sum_{j=k}^{T-1}\beta^{\,j-k}\chi_{k},$
which is equivalent to the right-hand side of (56). Therefore, (54) holds for $T\geq 2$.
From (54), we have
$\sum\limits_{t=0}^{T-1}$ ${\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}}=\sum\limits_{t=1}^{T-1}{\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}}$
$=\sum\limits_{j=0}^{T-2}{\sum\limits_{k=0}^{j}{{\beta^{j-k}}}{\chi_{k}}}=\sum\limits_{k=0}^{T-2}{\sum\limits_{j=k}^{T-2}{{\beta^{j-k}}}{\chi_{k}}}$
$=\sum\limits_{k=0}^{T-2}{{\chi_{k}}\sum\limits_{s=0}^{T-2-k}{{\beta^{s}}}}=\sum\limits_{k=0}^{T-2}{{\chi_{k}}}\frac{{1-{\beta^{T-1-k}}}}{{1-\beta}}$
$\leq\frac{1}{{1-\beta}}\sum\limits_{k=0}^{T-1}{{\chi_{k}}}.$ (57)
From (6.3) and (6.3), one has
$\sum\limits_{t=0}^{T-1}{{\chi_{t}}}\leq(max\{{\ell_{1}},{\ell_{2}}\}+\frac{{16{\ell_{3}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}})\sum\limits_{t=0}^{T-1}{{\chi_{t-1}}}$ (58)
$+\frac{{16{\ell_{3}}{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{{{(1-\beta)}^{2}}\gamma_{g}^{2}}}\sum\limits_{k=0}^{T-1}{{\chi_{k}}}+2T{\ell_{3}}{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}).$
From (LABEL:section6.2-chi_t_sum_simplified), one obtains
$(1$- $\frac{{16{\ell_{3}}{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{{{(1-\beta)}^{2}}\gamma_{g}^{2}}})\sum\limits_{t=0}^{T-1}{{\chi_{t}}}$
$\leq(max\{{\ell_{1}},{\ell_{2}}\}+\frac{{16{\ell_{3}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}})\sum\limits_{t=0}^{T-1}{{\chi_{t-1}}}$
$+2T{\ell_{3}}{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}),$ (59)
when ${\gamma_{g}}\geq\sqrt{\frac{{160{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}{\ell_{3}}{\eta^{2}}}}{{{{(1-\beta)}^{2}}}}}$, (6.3) can be rewritten as
$\sum\limits_{t=0}^{T-1}{{\chi_{t}}}\leq\rho_{1}\sum\limits_{t=0}^{T-1}{{\chi_{t-1}}}+\frac{{20T{\ell_{3}}{\eta^{2}}{\kappa_{4}}}}{9}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}).$ (60)
Then, we need to set the parameters such that ${\rho_{1}}<\frac{9}{{10}}$. First, note that ${\rho_{1}}<\frac{9}{{10}}$ holds if and only if ${\ell_{1}}+\rho{\ell_{3}}<\frac{9}{{10}}$ and ${\ell_{2}}+\rho{\ell_{3}}<\frac{9}{{10}}$.
For details, the parameter values in $\ell_{1}$$\ell_{3}$} are specified as
[…]
${\gamma_{x}}\in(max\{{m_{1}},{m_{3}}\},min\{{m_{2}},{m_{4}}\}),$
where
${{m_{1}}}=$ $\frac{{{a_{2}}+\sqrt{a_{2}^{2}-\frac{{24{a_{1}}\rho}}{\delta}}}}{{{-2{a_{1}}}}},$
${{m_{2}}}=$ $\sqrt[3]{{\frac{{12\rho{\omega^{3}}}}{{{\delta^{3}}{\omega^{3}}+{128\delta\omega{\lambda^{2}}+1500\rho\delta{\lambda^{2}}+25{\omega^{2}}\rho\delta{\lambda^{2}}}}}}},$
${m_{3}}=$ $\frac{{-{b_{2}}-\sqrt{b_{2}^{2}-\frac{{24{b_{1}}\rho}}{\delta}}}}{{2{b_{1}}}},$
${m_{4}}=$ $\min\{m_{4,1},m_{4,2}\},$
with ${a_{1}}=-\delta+\frac{{45\rho{\lambda^{2}}}}{{\omega\delta}}$, ${a_{2}}=-\frac{4}{{5}}+\rho-\frac{{\rho\omega}}{4}+\frac{{9\rho{\lambda^{2}}}}{{{\delta^{2}}}}$, ${b_{1}}=\frac{{3{\lambda^{2}}}}{\delta}+\frac{{45\rho{\lambda^{2}}}}{{\omega\delta}},\penalty 10000\ {b_{2}}={a_{2}}-\frac{\omega}{4}$, $m_{4,1}=\frac{{-{b_{2}}+\sqrt{b_{2}^{2}-\frac{{24{b_{1}}\rho}}{\delta}}}}{{2{b_{1}}}},\penalty 10000\ m_{4,2}=\sqrt[3]{{\frac{{12\rho{\omega^{3}}}}{{100\delta\omega{\lambda^{2}}+1500\delta\rho{\lambda^{2}}+25\delta{\omega^{2}}\rho{\lambda^{2}}}}}}\}$.
First, we substitute these values into $\ell_{1}$$\ell_{3}$}.
(i) For ${\varepsilon_{1}}(1+\alpha_{2}^{-1})$, due to ${\gamma_{x}}\delta\leq 1$, one has
${\varepsilon_{1}}(1+\alpha_{2}^{-1})={(1+\frac{{{\gamma_{x}}\delta}}{2})^{2}}{(1-{\gamma_{x}}\delta)^{2}}\leq{(1-\frac{{{\gamma_{x}}\delta}}{2})^{2}}.$
(ii) For $\kappa_{3}$}, based on $1+\frac{4}{\omega}\leq\frac{5}{\omega},1+\frac{\omega}{4}\leq\frac{5}{4}$, one has
${\kappa_{3}}=(1+\frac{4}{\omega})\gamma_{x}^{2}{\lambda^{2}}(1+\frac{\omega}{4})(1+\frac{4}{\omega})\leq\frac{{125}}{{4{\omega^{2}}}}\gamma_{x}^{2}{\lambda^{2}}\leq\frac{{32\gamma_{x}^{2}{\lambda^{2}}}}{{{\omega^{2}}}}.$
(iii) For $\varepsilon_{4}$}, based on ${(1+\frac{\omega}{4})^{3}}(1-\omega)\leq 1-\frac{\omega}{4}$, one has
${\varepsilon_{4}}$ $=(1+\frac{2}{{{\gamma_{x}}\delta}})\gamma_{x}^{2}{\lambda^{2}}(1+\frac{\omega}{4})(1+\frac{\omega}{4})(1-\omega)(1+\frac{\omega}{4})$
$\leq\frac{{3{\gamma_{x}}{\lambda^{2}}}}{\delta}(1-\frac{\omega}{4}).$
(iv) For $(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}})$, one has
$(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}})=(1+\frac{\omega}{4})(1-\omega)(1+\frac{\omega}{4})\leq 1-\frac{\omega}{4}.$
(v) For $\kappa_{2}$, due to ${(1+\frac{\omega}{4})^{3}}(1-\omega)\leq 1-\frac{\omega}{4}$, one has
${\kappa_{2}}$ $=(1+\frac{4}{\omega})\gamma_{x}^{2}{\lambda^{2}}(1+\frac{4}{\omega})(1+\frac{\omega}{4})(1+\frac{\omega}{4})(1-\omega)(1+\frac{\omega}{4})$
$\leq\frac{{25}}{{{\omega^{2}}}}\gamma_{x}^{2}{\lambda^{2}}(1-\frac{\omega}{4}).$
(vi) For $\varepsilon_{3}$}, based on the $(1+\frac{\omega}{4})(1-\omega)(1+\frac{4}{\omega})+(1+\frac{4}{\omega})\leq\frac{{10}}{\omega}$, $1+\frac{2}{{{\gamma_{x}}\delta}}\leq\frac{3}{{{\gamma_{x}}\delta}}$ and $1+\frac{{{\gamma_{x}}\delta}}{2}\leq\frac{3}{2}$, one has
${\varepsilon_{3}}$ $=(1+\frac{2}{{{\gamma_{x}}\delta}})\gamma_{x}^{2}{\lambda^{2}}(1+\frac{{{\gamma_{x}}\delta}}{2})[(1+\frac{\omega}{4})(1-\omega)(1+\frac{4}{\omega})$
$+(1+\frac{4}{\omega})]\leq\frac{{45{\gamma_{x}}{\lambda^{2}}}}{{\omega\delta}}.$
(vii) For $(1+{\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})$, based on $(1+\frac{{{\gamma_{x}}\delta}}{2}){(1-{\gamma_{x}}\delta)^{2}}\leq(1-\frac{{{\gamma_{x}}\delta}}{2})(1-{\gamma_{x}}\delta)\leq 1$, one has
$(1+$ ${\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})=(1+\frac{2}{{{\gamma_{x}}\delta}})(1+\frac{{{\gamma_{x}}\delta}}{2}){(1-{\gamma_{x}}\delta)^{2}}$
$+{(1+\frac{2}{{{\gamma_{x}}\delta}})^{2}}\gamma_{x}^{2}{\lambda^{2}}\leq\frac{3}{{{\gamma_{x}}\delta}}(1+\frac{{3{\gamma_{x}}{\lambda^{2}}}}{\delta}).$
(viii) For $\kappa_{1}$}, one has
${\kappa_{1}}$ $=(1+\frac{4}{\omega})\gamma_{x}^{2}{\lambda^{2}}\{(1+\frac{4}{\omega})(1+\frac{\omega}{4})[(1+\frac{\omega}{4})(1-\omega)$
$(1+\frac{4}{\omega})+(1+\frac{4}{\omega})]+(1+\frac{\omega}{4})\}\leq(\frac{{375}}{{{\omega^{3}}}}+\frac{{25}}{{4\omega}})\gamma_{x}^{2}{\lambda^{2}}.$
From (i)–(viii), we have
${\ell_{1}}$ $={\varepsilon_{1}}(1+\alpha_{2}^{-1})+{\kappa_{3}}\leq{(1-\frac{{{\gamma_{x}}\delta}}{2})^{2}}+\frac{{32\gamma_{x}^{2}{\lambda^{2}}}}{{{\omega^{2}}}};$ (61a)
${\ell_{2}}$ $={\varepsilon_{4}}+(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}})+{\kappa_{2}}$
$\leq(1-\frac{\omega}{4})(\frac{{3{\gamma_{x}}{\lambda^{2}}}}{\delta}+1+\frac{{25}}{{{\omega^{2}}}}\gamma_{x}^{2}{\lambda^{2}})$
$\leq 1-\frac{\omega}{4}+\frac{{3{\gamma_{x}}{\lambda^{2}}}}{\delta}+\frac{{25}}{{{\omega^{2}}}}\gamma_{x}^{2}{\lambda^{2}}$ (61b)
${\ell_{3}}$ $={\varepsilon_{3}}+(1+{\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})+(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}})+{\kappa_{1}}$
$\leq\frac{{45{\gamma_{x}}{\lambda^{2}}}}{{\omega\delta}}+\frac{3}{{{\gamma_{x}}\delta}}(1+\frac{{3{\gamma_{x}}{\lambda^{2}}}}{\delta})+1-\frac{\omega}{4}+(\frac{{375}}{{{\omega^{3}}}}+\frac{{25}}{{4\omega}})\gamma_{x}^{2}{\lambda^{2}}.$ (61c)
Now, we analyze ${\ell_{1}}+\rho{\ell_{3}}<\frac{9}{{10}}$. From (61a)–(61c), we have
${\ell_{1}}+\rho{\ell_{3}}=$ $(\frac{{{\delta^{2}}}}{4}+\frac{{32{\lambda^{2}}}}{{{\omega^{2}}}}+\frac{{375\rho{\lambda^{2}}}}{{{\omega^{3}}}}+\frac{{25\rho{\lambda^{2}}}}{{4\omega}})\gamma_{x}^{2}$
$+(-\delta+\frac{{45\rho{\lambda^{2}}}}{{\omega\delta}}){\gamma_{x}}+\frac{{3\rho}}{{{\gamma_{x}}\delta}}+\frac{1}{{10}}$
$+\rho-\frac{{\rho\omega}}{4}+\frac{{9\rho{\lambda^{2}}}}{{{\delta^{2}}}}<\frac{9}{{10}}.$
When ${\gamma_{x}}\leq m_{2}$ and ${\gamma_{g}}>\sqrt{\frac{{720{\lambda^{2}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\omega{\delta^{2}}}}}$ we have
${a_{1}}\gamma_{x}^{2}+{a_{2}}{\gamma_{x}}+\frac{{6\rho}}{\delta}<0,$ (62)
we can see that (62) holds when ${\gamma_{x}}>m_{1}$.
Similarly,
${\ell_{2}}+\rho{\ell_{3}}=$ $(\frac{{25{\lambda^{2}}}}{{{\omega^{2}}}}+\frac{{375\rho{\lambda^{2}}}}{{{\omega^{3}}}}+\frac{{25\rho{\lambda^{2}}}}{{4\omega}})\gamma_{x}^{2}$
$+(\frac{{3{\lambda^{2}}}}{\delta}+\frac{{45\rho{\lambda^{2}}}}{{\omega\delta}}){\gamma_{x}}+\frac{{3\rho}}{{{\gamma_{x}}\delta}}+\frac{1}{{10}}$
$-\frac{\omega}{4}+\frac{{9\rho{\lambda^{2}}}}{{{\delta^{2}}}}+\rho-\frac{{\rho\omega}}{4}<\frac{9}{{10}}.$
When ${\gamma_{x}}\leq m_{4,2}$, one has
${b_{1}}\gamma_{x}^{2}+{b_{2}}{\gamma_{x}}+\frac{{6\rho}}{\delta}<0,$ (63)
we can conclude that (63) holds when $\eta\leq\sqrt{\frac{{(16+5\omega)\gamma_{g}^{2}}}{{320{\varepsilon_{0}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}}$ and $m_{3}<{\gamma_{x}}<m_{4,1}$.
Building upon the above, we have
$max\{{\ell_{1}},{\ell_{2}}\}+\frac{{16{\ell_{3}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}<\frac{9}{{10}}.$
Then, we analyze that $\gamma_{x}$} is nonempty.
(i) We show an upper bound on $\max\{m_{1},m_{3}\}$. Using the above notations, $m_{1}$ can be rewritten as
${m_{1}}=\frac{{\sqrt{\varepsilon_{2}^{2}+{u_{1}}}-{\varepsilon_{2}}}}{{2{\varepsilon_{1}}}},$
where ${u_{1}}=\frac{{24{\varepsilon_{1}}\rho}}{\delta}$. Based on $\sqrt{\varepsilon_{2}^{2}+{u_{1}}}+{\varepsilon_{2}}\geq 2{\varepsilon_{2}}$, one obtains $\sqrt{\varepsilon_{2}^{2}+{u_{1}}}-{\varepsilon_{2}}=\frac{{{u_{1}}}}{{\sqrt{\varepsilon_{2}^{2}+{u_{1}}}+{\varepsilon_{2}}}}\leq\frac{{{u_{1}}}}{{2{\varepsilon_{2}}}}$. Then,
${m_{1}}\leq\frac{1}{{2{\varepsilon_{1}}}}\cdot\frac{{{u_{1}}}}{{2{\varepsilon_{2}}}}=\frac{{6\rho}}{{{\varepsilon_{2}}\delta}}.$
For ${{\gamma_{g}}}={{\tilde{\gamma}}_{1}}\leq\sqrt{40{\varepsilon_{0}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}$, one has
${\varepsilon_{2}}\geq\frac{4}{5}-\rho_{0}{\varepsilon_{0}}=\frac{2}{5},$
then,
$m_{1}\leq\frac{15}{\delta}\rho.$ (64)
Similarly, we can rewrite $m_{3}$ as
${m_{3}}=\frac{{{\varepsilon_{4}}-\sqrt{\varepsilon_{4}^{2}-{u_{2}}}}}{{2{b_{1}}}},$
where ${u_{2}}=\frac{{24{b_{1}}\rho}}{\delta}$. Based on ${\varepsilon_{4}}+\sqrt{\varepsilon_{4}^{2}-{u_{2}}}\geq{\varepsilon_{4}}$, we have ${\varepsilon_{4}}-\sqrt{\varepsilon_{4}^{2}-{u_{2}}}=\frac{{{u_{2}}}}{{{\varepsilon_{4}}+\sqrt{\varepsilon_{4}^{2}-{u_{2}}}}}\leq\frac{{{u_{2}}}}{{{\varepsilon_{4}}}}$. Then,
${m_{3}}\leq\frac{1}{{2{b_{1}}}}\cdot\frac{{{u_{2}}}}{{{\varepsilon_{4}}}},$
For $0<\rho\leq{\rho_{0}}$, one has
${\varepsilon_{4}}\geq\frac{2}{5}+\frac{\omega}{4}\geq\frac{2}{5},$
then,
${m_{3}}\leq\frac{{30\rho}}{\delta}.$ (65)
Combining (64) and (65) gives
$\max\{m_{1},m_{3}\}\leq\frac{30}{\delta}\rho.$ (66)
(ii) Next, we analyze $\min\{m_{2},m_{4}\}$.
A direct computation shows that ${\varepsilon_{9}}-{\kappa_{5}}=\delta^{3}\omega^{3}+28\delta\omega\lambda^{2}>0$. Recall that $m_{2}^{3}=\frac{12\rho\omega^{3}}{{\varepsilon_{9}}}$ and $m_{4,2}^{3}:=\frac{12\rho\omega^{3}}{{\kappa_{5}}}$. Hence,
$m_{2}<m_{4,2}.$ (67)
Using the previously defined $\varepsilon_{4}$, we obtain
${m_{4,1}}=\frac{{{\varepsilon_{4}}+\sqrt{\varepsilon_{4}^{2}-{u_{2}}}}}{{2{b_{1}}}}\geq\frac{{{\varepsilon_{4}}}}{{2{b_{1}}}}.$
For $0<\rho\leq\rho_{0}$, ${\varepsilon_{4}}(\rho)$ is increasing and $b_{1}(\rho)$ is decreasing. Therefore, $\varepsilon_{4}(\rho)\geq\varepsilon_{4}(\rho_{0}),b_{1}(\rho)\leq b_{1}(\rho_{0})$, then, $\frac{{{\varepsilon_{4}}(\rho)}}{{2{b_{1}}(\rho)}}\geq\frac{{{\varepsilon_{4}}({\rho_{0}})}}{{2{b_{1}}({\rho_{0}})}}$,
$m_{4,1}\geq\frac{{{\varepsilon_{4}}({\rho_{0}})}}{{2{b_{1}}({\rho_{0}})}}=\varepsilon_{5}.$
On the one hand, due to ${\varepsilon_{9}}\geq{\varepsilon_{6}}$, one has
$m_{2}^{3}=\frac{12\rho\omega^{3}}{{\varepsilon_{9}}}\leq\frac{12\rho\omega^{3}}{\varepsilon_{6}},$
i.e.,
$m_{2}\leq(\frac{12\omega^{3}}{\varepsilon_{6}})^{{\frac{1}{3}}}\rho^{\frac{1}{3}}.$
When $\gamma_{g}\leq{{\tilde{\gamma}}_{2}}=\sqrt{\frac{{192{\omega^{3}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\varepsilon_{5}^{3}{\varepsilon_{6}}}}}$, one has
$(\frac{12\omega^{3}}{\varepsilon_{6}})^{\frac{1}{3}}\rho^{\frac{1}{3}}\leq\varepsilon_{5},$
and
$m_{2}\leq\varepsilon_{5}\leq m_{4,1}.$ (68)
From (67) and (68), we have
$\min\{m_{2},m_{4}\}=m_{2}.$ (69)
On the other hand, using $\varepsilon_{9}$ is increasing in $\rho$ and $0<\rho\leq\rho_{0}$, we have $\varepsilon_{9}(\rho)\leq\varepsilon_{9}(\rho_{0})$. Hence,
$m_{2}^{3}=\frac{12\rho\omega^{3}}{\varepsilon_{9}(\rho)}\geq\frac{12\rho\omega^{3}}{\varepsilon_{7}},$
i.e.,
$m_{2}\geq{\varepsilon_{8}}\rho^{\frac{1}{3}}.$
Finally, when $\gamma_{g}\leq{{\tilde{\gamma}}_{3}}=\frac{{4\eta d{\sigma_{2}}{L_{{f_{1}}}}}}{{{{(\frac{{\delta{\varepsilon_{8}}}}{{30}})}^{\frac{3}{4}}}}}$, we have
$\frac{30}{\delta}\rho\leq{\varepsilon_{8}}\rho^{\frac{1}{3}}.$ (70)
From (66), (70), we conclude that
$\max\{m_{1},m_{3}\}\leq\frac{30}{\delta}\rho\leq{\varepsilon_{8}}\rho^{\frac{1}{3}}\leq\min\{m_{2},m_{4}\}.$
More notably, $\sum\limits_{t=0}^{T-1}{{\chi_{t-1}}}={\chi_{-1}}+\sum\limits_{t=0}^{T-2}{{\chi_{t}}}\leq{\chi_{-1}}+\sum\limits_{t=0}^{T-1}{{\chi_{t}}}$, one has
$\sum\limits_{t=0}^{T-1}{{\chi_{t}}}$ $\leq{\rho_{1}}({\chi_{-1}}+\sum\limits_{t=0}^{T-1}{{\chi_{t}}})+\frac{{20T{\ell_{3}}{\eta^{2}}{\kappa_{4}}}}{9}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}})$ (71)
$={\rho_{1}}\sum\limits_{t=0}^{T-1}{{\chi_{t}}}+{\rho_{1}}{\chi_{-1}}+\frac{{20T{\ell_{3}}{\eta^{2}}{\kappa_{4}}}}{9}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}).$
From (71), one obtains
$\sum\limits_{t=0}^{T-1}{{\chi_{t}}}$ $\leq\frac{{{\rho_{1}}{\chi_{-1}}}}{{1-{\rho_{1}}}}+\frac{{20T{\ell_{3}}{\eta^{2}}{\kappa_{4}}}}{{9(1-{\rho_{1}})}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}})$
$=\frac{{20T{\ell_{3}}{\eta^{2}}{\kappa_{4}}}}{{9(1-{\rho_{1}})}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}).$ (72)
By averaging over (6.3) over $t=0$ to t=T1t=T-1, we obtain
$\frac{1}{T}\sum\limits_{t=0}^{T-1}{{\chi_{t}}}$ $\leq\frac{{20{\ell_{3}}{\eta^{2}}{\kappa_{4}}}}{{9(1-{\rho_{1}})}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}})$
$\leq\frac{{25{\ell_{3}}{\eta^{2}}{\kappa_{4}}}}{{{{(1-\beta)}^{2}}}}.$ (73)
By virtue of $\eta=(1-\beta)\sqrt{\frac{n}{T}}$ and $\frac{1}{T}\sum\limits_{t=0}^{T-1}{\sum\limits_{i=1}^{n}{\mathbb{E}{{\|{{x_{i,t}}-{{\bar{x}}_{t}}}\|}^{2}}}}\leq\frac{1}{T}\sum\limits_{t=0}^{T-1}{{\chi_{t}}}$. Then, (6.3) can be rearranged as
$\frac{1}{T}\sum\limits_{t=0}^{T-1}$ ${\sum\limits_{i=1}^{n}{\mathbb{E}{{\|{{x_{i,t}}-{{\bar{x}}_{t}}}\|}^{2}}}}\leq\frac{{25{\ell_{3}}n{\kappa_{4}}}}{{{T}}}$
$\leq$ $\frac{{100{n^{2}}{\ell_{3}}{d^{2}}\sigma_{2}^{4}L_{{f_{1}}}^{2}}}{{{{T}}}}+\frac{{100{n^{2}}{\ell_{3}}{d^{2}}\sigma_{2}^{2}(\gamma_{1}^{2}+{\vartheta_{2}})}}{{{{T}}\gamma_{g}^{2}}}.$ (74)
Now, we show the upper bound of $\frac{1}{T}\sum\limits_{t=0}^{T-1}\|{\nabla f({{\bar{x}}_{t}})\|{{}^{2}}}$.
Based on the Assumption 1 and (6.2), one has
${f}({\bar{x}}_{t+1}^{a})$ $\leq f(\bar{x}_{t}^{a})+\langle{\nabla f(\bar{x}_{t}^{a}),\bar{x}_{t+1}^{a}-\bar{x}_{t}^{a}}\rangle+\frac{{{L_{{f_{2}}}}}}{2}{\|{\bar{x}_{t+1}^{a}-\bar{x}_{t}^{a}}\|^{2}}$ (75)
$\leq{f}({\bar{x}}_{t}^{a})-\langle{\nabla{f}({\bar{x}}_{t}^{a}),\frac{\eta}{{1-\beta}}{{\bar{g}}_{t}}}\rangle+\frac{{L_{{f_{2}}}}}{2}{(\frac{\eta}{{1-\beta}})^{2}}{\|{{{\bar{g}}_{t}}}\|^{2}},$
where ${{\bar{g}}_{t}}={\frac{1}{n}\sum\limits_{i=1}^{n}{{g_{i,t}}}}$. Based on the Lemma 1, we have
$\mathbb{E}[{{\bar{g}}_{t}}|{{\mathcal{H}}_{t}}]$ $=\frac{1}{n}\sum\limits_{i=1}^{n}{\mathbb{E}[{g_{i,t}}|{{\mathcal{H}}_{t}}]}$
$=d{\sigma_{1}}[\frac{1}{n}\sum\limits_{i=1}^{n}{(\nabla{F_{i}}({x_{i,t}})+{b_{i,t}})}]$
$=d{\sigma_{1}}[\nabla{{\bar{F}}_{i}}({x_{t}})+{{\bar{b}}_{t}}].$ (76)
Then,
$\|{b_{i,t}}\|$ $\leq\frac{{{\gamma_{g}}}}{{2{\sigma_{1}}}}\mathbb{E}{{}_{u}}[{{{\|{{u_{i,t}}}\|}_{2}}{{\|{{u_{i,t}}}\|}_{2}}{{\|{{\nabla^{2}}{F_{i}}(\upsilon)}\|}_{2}}{{\|{{u_{i,t}}}\|}_{2}}}$] (77)
$\leq\frac{{{\gamma_{g}}}}{{2{\sigma_{1}}}}\sigma_{2}^{3}{\vartheta_{1}},$
hence, $\|{\bar{b}_{t}}\|\leq\frac{{{\gamma_{g}}}}{{2{\sigma_{1}}}}\sigma_{2}^{3}{\vartheta_{1}}$. Then, we denote $\nabla{{\bar{F}}_{i}}({x_{t}})=\frac{1}{n}\sum\limits_{i=1}^{n}{\nabla{F_{i}}({x_{i,t}})}$, one has
$\|{\nabla{f}({\bar{x}}_{t}^{a})}$- ${\nabla{{\bar{F}}_{i}}({x_{t}})}\|=\|{\frac{1}{n}\sum\limits_{i=1}^{n}{(\nabla{F_{i}}({\bar{x}}_{t}^{a})-\nabla{F_{i}}({x_{i,t}}))}}\|$ (78)
$\leq$ $\frac{{L_{{f_{2}}}}}{n}\sum\limits_{i=1}^{n}{\|{{\bar{x}}_{t}^{a}-{x_{i,t}}}\|}\leq\frac{L_{{f_{2}}}}{{\sqrt{n}}}{({\sum\limits_{i=1}^{n}{{{\|{{\bar{x}}_{t}^{a}-{x_{i,t}}}\|}^{2}}}})^{\frac{1}{2}}},$
where the first inequality holds due to Assumption 2; the second inequality holds due to the Cauchy–Schwarz inequality.
Taking the conditional expectation given ${\mathcal{H}}_{t}$} on both sides of (75), one has
$\mathbb{E}[{{f}}$ ${({\bar{x}}_{t+1}^{a})|{{\mathcal{H}}_{t}}}]$
$\leq$ ${f}({\bar{x}}_{t}^{a})-\frac{\eta}{{1-\beta}}\mathbb{E}[{\langle{\nabla{f}({\bar{x}}_{t}^{a}),{{\bar{g}}_{t}}}\rangle}]+\frac{{L_{{f_{2}}}}}{2}{({\frac{\eta}{{1-\beta}}})^{2}}\mathbb{E}{{{\|{{{\bar{g}}_{t}}}\|}^{2}}}$
$\mathop{\leq}\limits^{(a)}$ ${f}({\bar{x}}_{t}^{a})-\frac{{\eta d{\sigma_{1}}}}{{1-\beta}}\mathbb{E}[{\langle{\nabla{f}({\bar{x}}_{t}^{a}),\nabla{{\bar{F}}_{i}}({x_{t}})+{{\bar{b}}_{t}}}\rangle}]$
$+\frac{{L_{{f_{2}}}}}{2}{({\frac{\eta}{{1-\beta}}})^{2}}(\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}\chi_{t}+\frac{{{\kappa_{4}}}}{n})$
$=$ ${f}({\bar{x}}_{t}^{a})-\frac{{\eta d{\sigma_{1}}}}{{1-\beta}}\langle\nabla f(\bar{x}_{t}^{a}),\nabla{{\bar{F}}_{i}}({x_{t}})+{{\bar{b}}_{t}}{+\nabla f(\bar{x}_{t}^{a})}$
$-{\nabla f(\bar{x}_{t}^{a})}\rangle+\frac{{L_{{f_{2}}}}}{2}{({\frac{\eta}{{1-\beta}}})^{2}}(\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}\chi_{t}+\frac{{{\kappa_{4}}}}{n})$
$\leq$ ${f}({\bar{x}}_{t}^{a})-\frac{{\eta d{\sigma_{1}}}}{{1-\beta}}{\|{\nabla f(\bar{x}_{t}^{a})}\|^{2}}-\frac{{\eta d{\sigma_{1}}}}{{1-\beta}}\langle{\nabla f(\bar{x}_{t}^{a}),{{\bar{b}}_{t}}}\rangle$
$+\frac{{\eta d{\sigma_{1}}}}{{1-\beta}}\langle{\nabla f(\bar{x}_{t}^{a}),\nabla f(\bar{x}_{t}^{a})-\nabla{{\bar{F}}_{i}}({x_{t}})}\rangle$
$+\frac{{L_{{f_{2}}}}}{2}{({\frac{\eta}{{1-\beta}}})^{2}}(\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}\chi_{t}+\frac{{{\kappa_{4}}}}{n})$
$\mathop{\leq}\limits^{(b)}$ ${f}({\bar{x}}_{t}^{a})-\frac{{\eta{d}{\sigma_{1}}}}{{1-\beta}}{\|{\nabla{f}({\bar{x}}_{t}^{a})}\|^{2}}+\frac{{\eta{d}{\sigma_{1}}}}{{1-\beta}}\|{\nabla{f}({\bar{x}}_{t}^{a})}\|\|{{{\bar{b}}_{t}}}\|\$
$+\frac{{\eta{d}{\sigma_{1}}}}{{1-\beta}}\|{\nabla{f}({\bar{x}}_{t}^{a})}\|\|{\nabla{f}({\bar{x}}_{t}^{a})-\nabla{{\bar{F}}_{i}}({x_{t}})}\|$ (79)
$+\frac{{L_{{f_{2}}}}}{2}{({\frac{\eta}{{1-\beta}}})^{2}}(\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}\chi_{t}+\frac{{{\kappa_{4}}}}{n})$
$\mathop{\leq}\limits^{(c)}$ ${f}({\bar{x}}_{t}^{a})-\frac{{\eta{d}{\sigma_{1}}}}{{1-\beta}}{\|{\nabla f(\bar{x}_{t}^{a})}\|^{2}}$
$+\frac{{L_{{f_{2}}}}}{2}{({\frac{\eta}{{1-\beta}}})^{2}}(\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}\chi_{t}+\frac{{{\kappa_{4}}}}{n})$
$+\frac{{\eta{d}{\sigma_{1}}}}{{1-\beta}}[\frac{1}{4}{\|{\nabla f(\bar{x}_{t}^{a})}\|^{2}}+{\|{{{\bar{b}}_{t}}}\|^{2}}]$
$+\frac{{\eta{d}{\sigma_{1}}}}{{1-\beta}}[\frac{1}{4}{\|{\nabla f(\bar{x}_{t}^{a})}\|^{2}}+{\|{\nabla f(\bar{x}_{t}^{a})-\nabla{{\bar{F}}_{i}}{{({x_{t}})}}}\|^{2}}]$
$=$ ${f}({\bar{x}}_{t}^{a})-\frac{{\eta{d}{\sigma_{1}}}}{{2(1-\beta)}}{\|{\nabla f(\bar{x}_{t}^{a})}\|^{2}}+\frac{{\eta{d}{\sigma_{1}}}}{{1-\beta}}{\|{{{\bar{b}}_{t}}}\|^{2}}$
$+\frac{{\eta{d}{\sigma_{1}}}}{{1-\beta}}{\|{\nabla f(\bar{x}_{t}^{a})-\nabla{{\bar{F}}_{i}}{{({x_{t}})}}}\|^{2}}$
$+\frac{{L_{{f_{2}}}}}{2}{({\frac{\eta}{{1-\beta}}})^{2}}(\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}\chi_{t}+\frac{{{\kappa_{4}}}}{n})$
$\mathop{\leq}\limits^{(d)}$ ${f}({\bar{x}}_{t}^{a})-\frac{{\eta{d}{\sigma_{1}}}}{{2(1-\beta)}}{\|{\nabla{f}({\bar{x}}_{t}^{a})}\|^{2}}+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}$
$+\frac{{L_{{f_{2}}}}}{2}{({\frac{\eta}{{1-\beta}}})^{2}}(\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}\chi_{t}+\frac{{{\kappa_{4}}}}{n})$
${\rm{+}}\frac{{\eta d{\sigma_{1}}}}{{1-\beta}}\frac{{{L_{f_{2}}^{2}}}}{n}\sum\limits_{i=1}^{n}{{{\|{{\bar{x}}_{t}^{a}-{x_{i,t}}}\|}^{2}}}$
$\mathop{\leq}\limits^{(e)}$ ${f}({\bar{x}}_{t}^{a})-\frac{{\eta d{\sigma_{1}}}}{{4(1-\beta)}}{\|{\nabla{f}({{\bar{x}}_{t}})}\|^{2}}+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}$
$+\frac{{4{L_{{f_{2}}}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}{{(1-\beta)}^{2}}}}\chi_{t}+\frac{{{\kappa_{4}}{L_{{f_{2}}}}{\eta^{2}}}}{{2n{{(1-\beta)}^{2}}}}$
$+\frac{{\eta d{{{{L_{f_{2}}^{2}}}}}{\sigma_{1}}}}{{2(1-\beta)}}{\|{{{\bar{x}}_{t}}-{\bar{x}}_{t}^{a}}\|^{2}}+\frac{{\eta d{\sigma_{1}}{{{L_{f_{2}}^{2}}}}}}{{{n}(1-\beta)}}\sum\limits_{i=1}^{n}{{{\|{{\bar{x}}_{t}^{a}-{x_{i,t}}}\|}^{2}}},$
where $(a)$ holds due to (6.3); $(b)$ and $(c)$ apply the Cauchy-Schwarz inequality. $(d)$ holds due to (77); $(e)$ holds due to ${\|{\nabla{f}({{\bar{x}}_{t}})}\|^{2}}\leq 2{\|{\nabla{f}({{\bar{x}}_{t}})-\nabla{f}({\bar{x}}_{t}^{a})}\|^{2}}+2{\|{\nabla{f}({\bar{x}}_{t}^{a})}\|^{2}}\leq 2{L_{f_{2}}^{2}}{\|{{{\bar{x}}_{t}}-{\bar{x}}_{t}^{a}}\|^{2}}+2{\|{\nabla{f}({\bar{x}}_{t}^{a})}\|^{2}}$. Rearranging this gives ${\|{\nabla{f}({\bar{x}}_{t}^{a})}\|^{2}}\geq\frac{1}{2}{\|{\nabla{f}({{\bar{x}}_{t}})}\|^{2}}-{L_{f_{2}}^{2}}{\|{{{\bar{x}}_{t}}-{\bar{x}}_{t}^{a}}\|^{2}}$.
From (79), we have
$\frac{{\eta d{\sigma_{1}}}}{{4(1-\beta)}}$ ${\|{\nabla f({{\bar{x}}_{t}})}\|^{2}}$
$\leq$ $f(\bar{x}_{t}^{a})-\mathbb{E}[f(\bar{x}_{t+1}^{a})|{{\mathcal{H}}_{t}}]+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}+\frac{{{\kappa_{4}}{L_{{f_{2}}}}{\eta^{2}}}}{{2n{{(1-\beta)}^{2}}}}$
$+\frac{{4{L_{{f_{2}}}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}{{(1-\beta)}^{2}}}}{\chi_{t}}+\frac{{\eta dL_{{f_{2}}}^{2}{\sigma_{1}}}}{{2(1-\beta)}}{\|{{{\bar{x}}_{t}}-\bar{x}_{t}^{a}}\|^{2}}$
$+\frac{{\eta d{\sigma_{1}}L_{{f_{2}}}^{2}}}{{n(1-\beta)}}[2\sum\limits_{i=1}^{n}{{{\|{{x_{i,t}}-{{\bar{x}}_{t}}}\|}^{2}}}+2\sum\limits_{i=1}^{n}{{{\|{{{\bar{x}}_{t}}-\bar{x}_{t}^{a}}\|}^{2}}}]$
$\leq$ $f(\bar{x}_{t}^{a})-\mathbb{E}[f(\bar{x}_{t+1}^{a})|{{\mathcal{H}}_{t}}]+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}$
$+\frac{{{\kappa_{4}}{L_{{f_{2}}}}{\eta^{2}}}}{{2n{{(1-\beta)}^{2}}}}+\frac{{5\eta dL_{{f_{2}}}^{2}{\sigma_{1}}}}{{2(1-\beta)}}{\|{{{\bar{x}}_{t}}-\bar{x}_{t}^{a}}\|^{2}}$
$+(\frac{{4{L_{{f_{2}}}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}{{(1-\beta)}^{2}}}}+\frac{{2\eta d{\sigma_{1}}L_{{f_{2}}}^{2}}}{{n(1-\beta)}}){\chi_{t}}.$ (80)
From (43) and $\theta_{t-1}:=\sum_{k=0}^{t-1}\beta^{t-1-k}=\frac{1-\beta^{t}}{1-\beta}$, we have
$\|\bar{x}_{t}-\bar{x}_{t}^{a}\|^{2}$ $=(\frac{\eta\beta^{2}}{1-\beta})^{2}\|\frac{1}{n}\sum_{i=1}^{n}m_{i,t-1}\|^{2}$
$=(\frac{\eta\beta^{2}}{1-\beta})^{2}\theta_{t-1}^{2}\|\sum_{k=0}^{t-1}\frac{\beta^{t-1-k}}{\theta_{t-1}}\frac{1}{n}\sum_{i=1}^{n}g_{i,k}\|^{2}$
$\leq(\frac{\eta\beta^{2}}{1-\beta})^{2}\theta_{t-1}^{2}\sum_{k=0}^{t-1}\frac{\beta^{t-1-k}}{\theta_{t-1}}\|\frac{1}{n}\sum_{i=1}^{n}g_{i,k}\|^{2}$
$=(\frac{\eta\beta^{2}}{1-\beta})^{2}\theta_{t-1}\sum_{k=0}^{t-1}{\beta^{t-1-k}}\|\frac{1}{n}\sum_{i=1}^{n}g_{i,k}\|^{2}$
$\leq\frac{\beta^{4}\eta^{2}}{(1-\beta)^{3}}\sum_{k=0}^{t-1}{\beta^{t-1-k}}\|\frac{1}{n}\sum_{i=1}^{n}g_{i,k}\|^{2}.$ (81)
From (80)–(6.3), we have
$\frac{{\eta d{\sigma_{1}}}}{{4(1-\beta)}}$ ${\|{\nabla f({{\bar{x}}_{t}})}\|^{2}}\leq f(\bar{x}_{t}^{a})-\mathbb{E}[f(\bar{x}_{t+1}^{a})|{\mathcal{H}_{t}}]+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}$
$+\frac{{{\kappa_{4}}{L_{{f_{2}}}}{\eta^{2}}}}{{2n{{(1-\beta)}^{2}}}}+(\frac{{4{L_{{f_{2}}}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}{{(1-\beta)}^{2}}}}+\frac{{2\eta d{\sigma_{1}}L_{{f_{2}}}^{2}}}{{n(1-\beta)}}){\chi_{t}}$
$+\frac{{5{\eta^{3}}dL_{{f_{2}}}^{2}{\sigma_{1}}{\beta^{4}}}}{{2{{(1-\beta)}^{4}}}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\|{\frac{1}{n}\sum\limits_{i=1}^{n}{{g_{i,k}}}}\|^{2}}.$ (82)
Then, from (6.3), taking expectation with respect to the entire past and average over $t=0$ to T1T-1 gives
$\frac{{\eta d{\sigma_{1}}}}{{4T(1-\beta)}}\sum\limits_{t=0}^{T-1}{\mathbb{E}\|{\nabla f({{\bar{x}}_{t}})}\|^{2}}$
$\leq$ $\frac{1}{T}\mathbb{E}[f({{\bar{x}}_{0}})-f({{\bar{x}}_{T}^{a}})]+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}+\frac{{{\kappa_{4}}{L_{{f_{2}}}}{\eta^{2}}}}{{2n{{(1-\beta)}^{2}}}}$
$+(\frac{{4{L_{{f_{2}}}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}{{(1-\beta)}^{2}}}}+\frac{{2\eta d{\sigma_{1}}L_{{f_{2}}}^{2}}}{{n(1-\beta)}})\frac{1}{T}\sum\limits_{t=0}^{T-1}{\chi_{t}}$
$+\frac{{5{\eta^{3}}dL_{{f_{2}}}^{2}{\sigma_{1}}{\beta^{4}}}}{{2{{(1-\beta)}^{4}}}}\frac{1}{T}\sum\limits_{t=0}^{T-1}{\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}(\frac{{8{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}}}{\chi_{t}}+\frac{{{\kappa_{4}}}}{n})}}$
$\leq$ $\frac{1}{T}\mathbb{E}[f({{\bar{x}}_{0}})-f({{\bar{x}}_{T}^{a}})]+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}+\frac{{{\kappa_{4}}{L_{{f_{2}}}}{\eta^{2}}}}{{2n{{(1-\beta)}^{2}}}}$
$+(\frac{{4{L_{{f_{2}}}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}{{(1-\beta)}^{2}}}}+\frac{{2\eta d{\sigma_{1}}L_{{f_{2}}}^{2}}}{{n(1-\beta)}}$
$+\frac{{20{\eta^{3}}{d^{3}}L_{{f_{2}}}^{2}{\sigma_{1}}{\beta^{4}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{{{(1-\beta)}^{5}}n\gamma_{g}^{2}}})\frac{1}{T}\sum\limits_{t=0}^{T-1}{{\chi_{t}}}+\frac{{5{\eta^{3}}dL_{{f_{2}}}^{2}{\sigma_{1}}{\beta^{4}}{\kappa_{4}}}}{{2n{{(1-\beta)}^{5}}}}.$ (83)
where the first inequality holds due to (48). If $\eta\leq min\{\frac{{{{(1-\beta)}^{3}}}}{{5d{L_{{f_{2}}}}{\sigma_{1}}{\beta^{4}}}},\frac{{n(1-\beta)}}{{2d{\sigma_{1}}L_{{f_{2}}}^{2}}}\}$, (6.3) can be rewritten as
$\frac{{\eta d{\sigma_{1}}}}{{4(1-\beta)}}\frac{1}{T}\sum\limits_{t=0}^{T-1}{{{\|{\nabla f({{\bar{x}}_{t}})}\|}^{2}}}$
$\leq$ $\frac{{f({{\bar{x}}_{0}})-f(\bar{x}_{T}^{a})}}{T}+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}$
$+(\frac{{8{L_{{f_{2}}}}{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{n\gamma_{g}^{2}{{(1-\beta)}^{2}}}}+1)\frac{1}{T}\sum\limits_{t=0}^{T-1}{{\chi_{t}}}+\frac{{{\kappa_{4}}{L_{{f_{2}}}}{\eta^{2}}}}{{n{{(1-\beta)}^{2}}}}$
$\leq$ $\frac{{f({{\bar{x}}_{0}})-f^{*}}}{T}+\frac{{\eta d\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{4{\sigma_{1}}(1-\beta)}}+\frac{{200{L_{{f_{2}}}}{\eta^{4}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}{\ell_{3}}{\kappa_{4}}}}{{n\gamma_{g}^{2}{{(1-\beta)}^{4}}}}$
$+\frac{{25{\ell_{3}}{\eta^{2}}{\kappa_{4}}}}{{{{(1-\beta)}^{2}}}}+\frac{{{\kappa_{4}}{L_{{f_{2}}}}{\eta^{2}}}}{{n{{(1-\beta)}^{2}}}},$ (84)
where the first inequality holds due to (42); the second inequality holds due to (6.3) and $f(\bar{x}_{T}^{a})\geq f^{*}$ with $f^{*}$ is the minimum function value of the problem (1).
From (6.3), we have
$\frac{1}{T}$ $\sum\limits_{t=0}^{T-1}{{{\|{\nabla f({{\bar{x}}_{t}})}\|}^{2}}}$
$\leq$ $\frac{{4(1-\beta)(f({{\bar{x}}_{0}})-{f^{*}})}}{{\eta Td{\sigma_{1}}}}+\frac{{\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{\sigma_{1}^{2}}}+\frac{{4{\kappa_{4}}{L_{{f_{2}}}}\eta}}{{nd{\sigma_{1}}(1-\beta)}}$
$+\frac{{800{L_{{f_{2}}}}{\eta^{3}}d\sigma_{2}^{2}L_{{f_{1}}}^{2}{\ell_{3}}{\kappa_{4}}}}{{n\gamma_{g}^{2}{\sigma_{1}}{{(1-\beta)}^{3}}}}+\frac{{100{\ell_{3}}\eta{\kappa_{4}}}}{{d{\sigma_{1}}(1-\beta)}}$
$=$ $\frac{{4(1-\beta)(f({{\bar{x}}_{0}})-f(\bar{x}_{T}^{a}))}}{{\eta Td{\sigma_{1}}}}+\frac{{\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{\sigma_{1}^{2}}}+\eta\sigma_{2}^{2}L_{{f_{1}}}^{2}({d_{1}}+{d_{3}})$
$+\frac{{\eta(\gamma_{1}^{2}+{\vartheta_{2}})}}{{\gamma_{g}^{2}}}({d_{1}}+{d_{3}})+\frac{{\sigma_{2}^{2}L_{{f_{1}}}^{2}{\eta^{3}}{d_{2}}}}{{\gamma_{g}^{2}}}+\frac{{(\gamma_{1}^{2}+{\vartheta_{2}}){\eta^{3}}{d_{2}}}}{{\gamma_{g}^{4}}},$ (85)
where
${d_{1}}$ $=\frac{{16{L_{{f_{2}}}}d\sigma_{2}^{2}}}{{{\sigma_{1}}(1-\beta)}},$
${d_{2}}$ $=\frac{{3200{L_{{f_{2}}}}{d^{3}}\sigma_{2}^{4}L_{{f_{1}}}^{2}{\ell_{3}}}}{{{\sigma_{1}}{{(1-\beta)}^{3}}}},$
${d_{3}}$ $=\frac{{400nd\sigma_{2}^{2}{\ell_{3}}}}{{{\sigma_{1}}(1-\beta)}}.$
By choosing $\eta=(1-\beta)\sqrt{\frac{n}{T}}$ and running the algorithm for $T\geq\max\{{c_{1}},{c_{2}},{c_{3}},{c_{4}},{c_{5}},{c_{6}},{c_{7}}\}$ iterations, where
${c_{1}}$ $=\frac{{5nd{L_{{f_{2}}}}{\sigma_{1}}{\beta^{4}}}}{{1-\beta}},$
${c_{2}}$ $=\frac{{4{d^{2}}\sigma_{1}^{2}L_{{f_{2}}}^{4}}}{n},$
${c_{3}}$ $=\frac{{320n{{(1-\beta)}^{2}}{\varepsilon_{0}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(16+5\omega)\gamma_{g}^{2}}},$
${c_{4}}$ $=\frac{{720n{\lambda^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}{{(1-\beta)}^{2}}}}{{\gamma_{g}^{2}\omega{\delta^{2}}}},$
${c_{5}}$ $=\frac{{40n{{(1-\beta)}^{2}}{\varepsilon_{0}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}},$
${c_{6}}$ $=\frac{{192{\omega^{3}}n{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}(1-\beta)}}{{\gamma_{g}^{2}\varepsilon_{5}^{3}{\varepsilon_{6}}}},$
${c_{7}}$ $=\frac{{16n{{(1-\beta)}^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}{{(\frac{{\delta{\varepsilon_{8}}}}{{30}})}^{\frac{3}{2}}}}}.$
Then, (6.3) can be further simplified as
$\frac{1}{T}$ $\sum\limits_{t=0}^{T-1}{{{\|{\nabla f({{\bar{x}}_{t}})}\|}^{2}}}$
$\leq$ $\frac{{4(f({{\bar{x}}_{0}})-{f^{*}})}}{{\sqrt{nT}d{\sigma_{1}}}}+\frac{{\gamma_{g}^{2}\sigma_{2}^{6}\vartheta_{1}^{2}}}{{\sigma_{1}^{2}}}+\frac{{(1-\beta)\sqrt{n}\sigma_{2}^{2}L_{{f_{1}}}^{2}({d_{1}}+{d_{3}})}}{{\sqrt{T}}}$
$+\frac{{(1-\beta)\sqrt{n}(\gamma_{1}^{2}+{\vartheta_{2}})}}{{\sqrt{T}\gamma_{g}^{2}}}({d_{1}}+{d_{3}})+\frac{{\sigma_{2}^{2}L_{{f_{1}}}^{2}\sqrt{{n^{3}}}{{(1-\beta)}^{3}}{d_{2}}}}{{\sqrt{{T^{3}}}\gamma_{g}^{2}}}$
$+\frac{{(\gamma_{1}^{2}+{\vartheta_{2}})\sqrt{{n^{3}}}{{(1-\beta)}^{3}}{d_{2}}}}{{\sqrt{{T^{3}}}\gamma_{g}^{4}}}.$ (86)
This completes the proof.