### v2

# Compressed Momentum-based Single-Point Zero-Order Algorithm for Stochastic Distributed Nonconvex Optimization

Abstract:
This paper studies a compressed momentum-based single-point zeroth-order algorithm for stochastic distributed nonconvex optimization, aiming to alleviate communication overhead and address the unavailability of explicit gradient information. In the developed framework, each agent has access only to stochastic zeroth-order information of its local objective function, performs local stochastic updates with momentum, and exchanges compressed updates with its neighbors. We theoretically prove that the proposed algorithm achieves the convergence rate $\mathcal{O}(\frac{1}{\sqrt[4]{T}})$ to the exact solution with diminishing step sizes and can achieve a sublinear convergence rate ${\mathcal{O}}(\frac{1}{{\sqrt{T}}})$ towards a neighborhood of the stationary point with fixed step sizes. Numerical experiments validate the effectiveness and communication efficiency of the proposed algorithm.
keywords:
Compressed communication, distributed nonconvex optimization, momentum-based method, zeroth-order algorithm
1 Introduction
With the rapid growth of datasets and the expansion of network scales, distributed optimization has attracted significant attention due to its enhanced parallel computational capabilities and inherent robustness, where agents try to cooperatively solve a problem with information exchange only limited to immediate neighbors in the network. In many networked systems, such as power systems (Mao et al. (2019)) and wireless sensor networks (Kuthadi et al. (2022), Lei et al. (2022)), each computational node is typically limited to accessing information from its immediate neighbors, driven by requirements for security and scalability.
To date, a wide variety of distributed algorithms have been proposed, including distributed gradient tracking algorithms (Lei et al. (2022), Chen et al. (2024)), distributed ADMM (Carnevale et al. (2025), Gholami et al. (2023)), distributed pull-push gradient-based algorithm (Nedić et al. (2025), Dimlioglu and Choromanska (2025)) and so on. Additionally, a popular SGD heuristic that has proven to be crucial in many applications is the use of momentum, i.e., the use of a weighted average of past gradients instead of the current gradient (Sutskever et al. (2013)). Subsequently, Xia et al. (2024) investigated a paradigm of momentum-based systems for nonconvex constrained optimization and distributed nonconvex optimization. Li et al. (2025) proposed a gradient tracking stochastic distributed optimization algorithm with adaptive momentum.
However, in fields such as biochemistry, commercial applications and complex engineering systems, accessing gradients may pose significant challenges due to data privacy concerns, black-box constraints, or the high computational costs involved. To handle this problem, the gradient is approximated through random sampling and finite differences in zeroth-order (ZO) optimization. Existing ZO optimization algorithms can be divided into the following categories, namely, ZO with one-point estimator and ZO with two- and multiple-point estimator. Flaxman et al. (2004) proposed the ZO algorithm with one-point estimator that queries one function value at each iteration to estimate the gradient. Nowadays, some research findings are available to integrate ZO optimization with momentum-based algorithms. For instance, Huang et al. (2022) studied a class of accelerated zeroth-order and first-order momentum methods for both nonconvex mini-optimization and minimax-optimization. Qian and Zhao (2023) explored a ZO proximal stochastic recursive momentum algorithm.
Note that the distributed optimization algorithms require each agent to exchange information with its neighboring agents to obtain the global information. However, the communication channel often has limited bandwidth. Hence, a series of compression techniques have been developed to reduce communication overhead, including standard uniform quantizers (Xu et al. (2024)), unbiased compressors (Condat et al. (2022)), and sparsification methods (Wangni et al. (2018)).
In this paper, motivated by the well-known momentum-based optimization algorithm and by the practical limitation that querying multiple function values is often infeasible, especially in stochastic environments, we develop a Compressed Momentum-based Single-Point Zeroth-Order (CMSPZO) algorithm for stochastic distributed nonconvex optimization. The contributions are summarized as follows.
  • We propose a momentum-based method for stochastic distributed nonconvex optimization. Unlike Singh et al. (2021), we consider a more challenging gradient-free distributed setting, where gradients are approximated via noisy single-point function evaluations. Compared with Mhanna and Assaad (2023), our method incorporates compressed communication, which substantially reduces the communication overhead in distributed networks.
  • We establish that the proposed CMSPZO algorithm can achieve a convergence rate $\mathcal{O}(\frac{1}{\sqrt[4]{T}})$ towards the exact solution with diminishing step sizes and can achieve the sublinear convergence rate ${\mathcal{O}}(\frac{1}{{\sqrt{T}}})$ towards a neighborhood of a stationary point with fixed step sizes, where $T$ is the number of iterations.
  • In contrast to Mhanna and Assaad (2023), which assumes that $\|x\|$ is uniformly bounded. In this paper, we analyze two related auxiliary components to avoid the need for uniform boundedness assumption on $\|x\|$.
NotationsNotations: Let $\mathbb{R}$ denote the set of real numbers. Vectors with all entries equal to 11 or 0 are denoted by $\mathbf{1}$ and $\mathbf{0}$, respectively. $\|\cdot\|$ and $\|\cdot\|_{F}$ symbolize the $\ell_{2}$–norm and the Frobenius norm, respectively. The set of $n$-dimensional real vectors is denoted by $\mathbb{R}^{n}$, and $\mathbf{I}$ means the identity matrix. For a differentiable function $f$, $\nabla f$ represents its gradient. For two vectors $\in\mathbb{R}^{n}, a,b\langle a,b\rangle$ is the standard inner product. For a positive semi-definite matrix $\mathbf{W}$, $\lambda(\mathbf{W})$ denotes its spectral radius. Additionally, $\operatorname{sgn}(\cdot)$, $\circ$, $\lfloor\cdot\rfloor$, and $|\cdot|$ denote the element–wise sign, Hadamard product, floor function, and absolute value, respectively.
2 Preliminaries and problem formulation
2.1 Problem formulation
Consider a set of agents ${\mathcal{V}}=\{1,2,...,n\}$ connected by a communication network. Each agent $i$ is associated with a local objective function $F_{i}:{\mathbb{R}^{d}}\mapsto\mathbb{R}$. All agents aim to collaboratively find the decision variable $x\in{\mathbb{R}^{d}}$} that minimizes the stochastic global objective function:
$\mathop{\min}\limits_{x\in{\mathbb{R}^{d}}}{f}(x):=\frac{1}{n}\sum\limits_{i=1}^{n}{{F_{i}}(x)},$ (1)
where
${F_{i}}(x)\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}{\mathbb{E}_{\xi_{i}\sim{\mathcal{D}}_{i}}}[{f_{i}}(x,{\xi_{i}})],$ (2)
with $x$ denotes the global decision vector and $\xi_{i}$} is the local data that follows the local distribution ${\mathcal{D}}_{i}$, and ${f_{i}}(x,{\xi_{i}})$ is the stochastic local cost function of each agent ii. In addition, we assume that $\mathbb{E}_{\xi_{i}\sim{\mathcal{D}}_{i}}{{{[{{f_{i}}({x},{\xi_{i}})}]}^{2}}}\leq\gamma_{1}^{2}$, where $\gamma_{1}>0$.
Assumption 1 (The local functions)
We assume the Lipschitz continuity of the local objective functions $f_{i}(x,\xi_{i})$ such that
$\|f_{i}(x,\xi_{i})-f_{i}(x^{\prime},\xi_{i})\|\leq L_{f_{1}}\|x-x^{\prime}\|,\quad\forall x,x^{\prime}\in\mathbb{R}^{d},$ (3)
where $L_{f_{1}}$ denotes the Lipschitz constant.
Assumption 2 (The objective function)
Both $\nabla F_{i}(x)$ and $\nabla^{2}F_{i}(x)$ exist and are continuous, and there exists a constant $L_{{f_{2}}}>0$ such that $\|\nabla^{2}F_{i}(x)\|\leq L_{{f_{2}}}$, $\forall x\in\mathbb{R}^{d}.$ Hence, we can say that the stochastic objective function $f(x)$ is $L_{{f_{2}}}$–smooth:
$f(x)\leq f(x^{\prime})+\langle\nabla f(x^{\prime}),x-x^{\prime}\rangle+\frac{L_{{f_{2}}}}{2}\|x-x^{\prime}\|^{2}.$ (4)
To address the problem (1), we focus on the following aspects: exchanging information with other agents and collecting information about the local cost function. We introduce the following mechanisms to address these two aspects, together with several additional assumptions.
2.2 Graph Theory
An undirected graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ is introduced to describe the communication network with $n$ agents, where $\mathcal{V}=\{1,...,n\}$ is the set of agents. Define ${\mathcal{E}}\subset{\mathcal{V}}\times{\mathcal{V}}$ as the set of communication channels. The set of neighbors of the agent is denoted by ${{\mathcal{N}}_{i}}=\{j|(i,j)\in{\mathcal{E}}\}$. The non-negative weighted matrix $\mathbf{W}$ is represented as $\mathbf{W}=[{w_{ij}}]\in{{\mathbb{R}}^{n\times n}}$. And all diagonal elements $w_{ii}$ are strictly positive. An channel $(i,j)\in{\mathcal{E}}$ indicates that the agent $i$ can communicate information with the agent $j$. Furthermore, ${w_{ij}}>0$ if $(i,j)\in{\mathcal{E}}$ and otherwise ${w_{ij}}=0$.
Assumption 3
Assume that the graph $\mathcal{G}$} is connected. The weighted matrix $\mathbf{W}$ is symmetric and doubly stochastic, i.e., ${\mathbf{1}^{T}}\mathbf{W}={\mathbf{1}^{T}},\mathbf{W}\mathbf{1}=\mathbf{1}$.
As shown in Assumption 3, for such a $\mathbf{W}$ associated with $\mathcal{G}$}, it is known that $\lambda_{1}(\mathbf{W})=1$ and $\lambda_{i}(\mathbf{W})\in(-1,1)$ for all $i=2,\dots,n$. The spectral gap $\delta\in(0,1]$ is defined as $\delta:=1-|\lambda_{2}(\mathbf{W})|$.
2.3 Zeroth-order optimization
This section concentrates on utilizing the stochastic zeroth-order (ZO) information to estimate the gradients of local objective functions. In the absence of explicit gradient expressions, we assume each agent could query the local function value at one point.
Let $F(x)=\mathbb{E}_{\xi}[f(x,\xi)]$, where $f(x,\xi)$ denotes a stochastic differentiable function. As shown in Flaxman et al. (2004), the one-point estimator can be constructed as
$g(x,\xi):=\frac{d}{\gamma}f(x+\gamma u,\xi)u,$ (5)
where $x$ means the optimization variable, $\gamma>0$ denotes the parameter that sets the level of smoothing. And $u$ is a random vector that follows a symmetric distribution.
To solve the problem (1), agent $i$ can only query the function values of $f_{i}$ at exactly one point. These function queries are assumed as $\tilde{f}_{i}=f_{i}+\varphi_{i}$. We note that $\varphi_{i}$ is added to differentiate between the stochastic nature of $f_{i}$ and the noise imposed on its querying.
Then, we assume that the function queries are
$\tilde{f}_{i}({x_{i,t}}+{\gamma_{g}}{u_{i,t}},{\xi_{i,t}})=f_{i}({x_{i,t}}+{\gamma_{g}}{u_{i,t}},{\xi_{i,t}})+\varphi_{i,t},$
where $x_{i,t}\in\mathbb{R}^{d}$ is the estimation of $x$ maintained by agent $i$ at time step $t$; $\varphi_{i,t}$ denotes the additive noise at time step $t$; $\gamma_{g}$ is a positive parameter called the smoothing radius; $u_{i,t}$ is a random perturbation vector generated by the agent $i$ at time step $t$.
Inspired by (5), we have
$g_{i,t}$ $=\frac{d}{\gamma_{g}}u_{i,t}\tilde{f}_{i}\big(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t}\big)$
$=\frac{d}{\gamma_{g}}u_{i,t}(f_{i}(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t})+\varphi_{i,t}).$ (6)
For the noise term $\varphi_{i,t}$ and the perturbation vector $u_{i,t}$, we adopt the following assumptions.
Assumption 4
The noise term $\varphi_{i,t}$ is assumed to have a zero mean and be uncorrelated with bounded variance, where $\mathbb{E}(\varphi_{i,t})=0$ and $\mathbb{E}(\varphi_{i,t}^{2})=\vartheta_{2}<\infty,i\in\mathcal{N}$, and $\mathbb{E}(\varphi_{i,t}\varphi_{j,t})=0$ for $i\neq j$.
Assumption 5
The perturbation vector $u_{i,t}=(u_{i,t,1},\ldots,\\ u_{i,t,d})^{\top}\in{\mathbb{R}^{d}}$ is chosen independently by each agent $i\in\mathcal{N}$ from others and previous samples. In addition, the elements of $u_{i,t}$ are assumed i.i.d. with $\mathbb{E}(u_{i,t,d})^{2}=\sigma_{1}>0$, which implies that $\mathbb{E}[u_{i,t}u_{i,t}^{T}]=\sigma_{1}I_{d}$. Then, we assume that there exists a constant $\sigma_{2}>0$ such that $\|u_{i,t}\|\leq\sigma_{2}.$
Remark 1
An example of a perturbation vector satisfying Assumption 5, we can choose every dimension of $u_{i,t}$ from the symmetric Bernoulli distribution on $\{-\tfrac{1}{\sqrt{d}},\tfrac{1}{\sqrt{d}}\}$. Then, $\sigma_{1}=\frac{1}{d},\sigma_{2}=1$.
Lemma 1
(Proposition $3.$ of Mhanna and Assaad (2023)) Under Assumptions 4 and 5, then, $g_{i,t}$ is a biased estimator of the agent’s gradient $\nabla{F_{i}}({x_{i,t}}),\forall i\in{\mathcal{V}}$ for every $t\geq 0$, i.e.,
$\mathbb{E}{{}_{u,\xi,\varphi}}[{g_{i,t}}|{{\mathcal{H}}_{t}}]=d{\sigma_{1}}[\nabla{F_{i}}({x_{i,t}})+{b_{i,t}}],$ (7)
with
${{b_{i,t}}}={\frac{{{\gamma_{g}}}}{{2{\sigma_{1}}}}\mathbb{E}{{}_{u}}[{u_{i,t}}u_{i,t}^{T}{\nabla^{2}}{F_{i}}(\upsilon){u_{i,t}}|{{\mathcal{H}}_{t}}}$], (8)
where ${{b_{i,t}}}\in\mathbb{R}^{d}$ is the bias with respect to the true gradient, and $\upsilon\in[x_{i,t},\,x_{i,t}+\gamma_{g}u_{i,t}]$. Moreover, ${{\mathcal{H}}_{t}}=\{{{\mathbf{x}_{0}}},{\xi_{0}},...,{{\mathbf{x}_{t-1}}},{\xi_{t-1}},{{\mathbf{x}_{t}}}\}$ denotes the history sequence, where $\mathbf{x}_{t}$}} is defined in Section 3.2.
2.4 Compression operators
To reduce communication overhead, we consider a scenario where information exchange between agents is compressed. Specifically, we consider the compressors $\mathcal{C}(\cdot)$ with bounded relative compression error satisfying the following assumption.
Definition 1 (Compression)
A function $\mathcal{C}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is called a compression operator, if there exists a positive constant $\omega\in(0,1]$, such that for every $x\in\mathbb{R}^{d}$:
$\mathbb{E}_{\mathcal{C}}[\|x-\mathcal{C}(x)\|_{2}^{2}]\leq(1-\omega)\|x\|_{2}^{2},$ (9)
where expectation is taken over the randomness of $\mathcal{C}$. We assume that $\mathcal{C}(0)=0$.
Remark 2
As stated in Liao et al. (2022) is naturally satisfied by various compressors in distributed algorithms. For example, the Randk sparsifier (Beznosikov et al. (2023)), the Topk sparsifier (Zou et al. (2022)) and norm-sign compressor (Yi et al. (2022)).
3 Main results
In this section, we propose a Compressed Momentum-based Single-Point Zeroth-Order (CMSPZO) algorithm. First, we introduce the CMSPZO algorithm and its motivation in Section 3.1. Then, we provide the supporting lemmas and establish the convergence of the CMSPZO algorithm in Section 3.2.
3.1 Algorithm Description
When the gradient information is available, in Yu et al. (2019), the following distributed first-order momentum-based algorithm was proposed to solve (1):
$m_{i,t}$ $=\beta m_{i,{t-1}}+\nabla F_{i}(x_{i,t-1},\xi_{i,t-1}),$ (10a)
$x_{i,{t}}$ $=x_{i,{t-1}}-\eta m_{i,t},$ (10b)
where $m_{i,t}$ is the local momentum of agent $i$ at time step $t$; $\eta$ means stochastic gradient descent (SGD) step-size; $\beta\in[0,1)$ denotes momentum coefficient.
More notably, the first-order information is available, which may not always hold in practice, as gradient computation can be impractical or too costly for large-scale or highly complex systems. Then, we replace the gradient in (10a) with the stochastic gradient estimator (6).
To realize exchange of compressed parameters between agents, each agent $i$ maintains an estimate $\hat{x}_{i,t}$ of $x_{i,t}$ at time step $t$. Specifically, the updates for agent $i\in\mathcal{V}$ can be formulated as follows:
${\hat{x}_{j,t+1}}={\hat{x}_{j,t}}+{\mathcal{C}}({x_{i,t+\frac{1}{2}}}-{\hat{x}_{i,t}}),$ (11)
with
${x_{i,t+\frac{1}{2}}}={x_{i,t}}-\eta(\beta{m_{i,t}}+{g_{i,t}}),$ (12)
where $x_{i,t+\frac{1}{2}}$} is formed by every agent computes a stochastic gradient $g_{i,t}$ and takes a local SGD step on $x_{i,t}$ using momentum-based method.
Based on (6), (10a), (11) and (12), we summarize the algorithm updates in Algorithm 1.
Algorithm 1 CMSPZO Algorithm
1:Input: Stopping time $T$, adjacency matrix $\mathbf{W}$, and positive parameters $\gamma_{x}$, $\eta$, $\beta$, $\gamma_{g}$.
2:Initialization: For all $i\in\mathcal{V}$, initial variable $x_{i,0}\in\mathbb{R}^{d}$, ${x}_{i,-1}:=\textbf{0}$, $\hat{x}_{i,-1}:=\textbf{0}$, $m_{i,-1}=\textbf{0}\in\mathbb{R}^{d}$.
3:for $t=0$ to T1T-1 in parallel on all agents $i$ do
4:  One-point zeroth-order gradient:
  • Sample $\xi_{i,t}$ from the distribution $\mathcal{D}_{i}$;
  • Sample $u_{i,t}$ from the symmetrical distribution;
  • Query $f_{i}(x_{i,t}+\gamma_{g}u_{i,t},\xi_{i,t})$;
  • Compute stochastic gradient $g_{i,t}$ using (6).
5:  Update $m_{i,t}$ by
${m_{i,t}}=\beta{m_{i,t-1}}+{g_{i,t}}.$ (13)
6:  Update $x_{i,t+\frac{1}{2}}$ according to (12).
7:  Compute $C(x_{i,t+\frac{1}{2}}-\hat{x}_{i,t})$ and broadcast it to its neighbors $\mathcal{N}_{i}$ according to (11).
8:  Consensus:
${x_{i,t+1}}={x_{i,t+\frac{1}{2}}}+{\gamma_{x}}\sum\limits_{j\in{{\mathcal{V}}_{i}}}{{w_{ij}}}({\hat{x}_{j,t+1}}-{\hat{x}_{i,t+1}}).$ (14)
9:end for
10:Output: $\{x_{i}(T)\}$.
3.2 Supporting Lemmas
For the sake of analysis, we introduce the following notations:
${{\mathbf{x}_{t}}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{x_{1,t}^{T}},.,{x_{n,t}^{T}}]^{T},\penalty 10000\ \penalty 10000\ {{\bar{\mathbf{x}}}_{t}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{{\bar{x}}_{t}^{T}},.,{{\bar{x}}_{t}^{T}}]^{T},$
${{\hat{\mathbf{x}}_{t}}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{{\hat{x}}_{1,t}^{T}},.,{{\hat{x}}_{n,t}^{T}}]^{T},\penalty 10000\ \penalty 10000\ {\mathbf{m}_{t}}\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{m_{1,t}^{T}},.,{m_{n,t}^{T}}]^{T},$
${\mathbf{g}_{t}}$ $\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[{g_{1,t}^{T}},.,{g_{n,t}^{T}}]^{T},\penalty 10000\ \penalty 10000\ \mathcal{C}(\mathbf{x})\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}[\mathcal{C}(x_{1})^{T},.,\mathcal{C}(x_{n})^{T}]^{T},$
where ${\bar{x}}_{t}=\frac{1}{n}\sum\limits_{i=1}^{n}{{x_{i,t}}}$ denotes the average of node parameters.
To begin with, we consider the following Lyapunov candidate function
${\chi_{t}}=\mathbb{E}\|{{\mathbf{x}_{t}}-{{\bar{\mathbf{x}}}_{t}}}\|_{F}^{2}+\mathbb{E}\|{{\mathbf{x}_{t}}-{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2},$ (15)
where incorporates the following error terms: the consensus error term $\mathbb{E}\|{{\mathbf{x}_{t}}-{{\bar{\mathbf{x}}}_{t}}}\|_{F}^{2}$ and the compressed error term $\mathbb{E}\|{{\mathbf{x}_{t}}-{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2}$. They play the key role for obtaining the convergence results of Algorithm 1.
Lemma 2
(Consensus error) Let $\{\mathbf{x}_{t}\}$ be the sequence generated by Algorithm 1, we have the following bound
$\mathbb{E}$ $\|\mathbf{x}_{t}-\bar{\mathbf{x}}_{t}\|_{F}^{2}$ (16)
$\leq$ ${\varepsilon_{1}}(1+\alpha_{2}^{-1})\mathbb{E}\|{{{\bar{\mathbf{x}}}_{t-1}}-{{\mathbf{x}}_{t-1}}}\|_{F}^{2}+{{{{\varepsilon_{4}}\mathbb{E}}}{\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}}}$
$+[(1+{\alpha_{2}})({\varepsilon_{1}}+{\varepsilon_{2}})+{\varepsilon_{3}}]\mathbb{E}\|{\eta(\beta{{\mathbf{m}}_{t-1}}+{{\mathbf{g}}_{t-1}})}\|_{F}^{2}.$
where $\alpha_{2}>0$ and $\varepsilon_{1}$$\varepsilon_{4}$} are given in Appendix 6.2.
Proof. See Appendix 6.2 for the proof.
Lemma 3
(Compressed error) Let $\{\mathbf{x}_{t}\}$ be the sequence generated by Algorithm 1, we have the following bound
$\mathbb{E}\|{{{\mathbf{x}}_{t}}-{{\hat{\mathbf{x}}}_{t}}}\|_{F}^{2}\leq{\kappa_{3}}\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\bar{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$ (17)
$+\{(1+{\alpha_{5}})(1-\omega)(1+{\alpha_{8}})+{\kappa_{2}}\}\mathbb{E}\|{{{\mathbf{x}}_{t-1}}-{{\hat{\mathbf{x}}}_{t-1}}}\|_{F}^{2}$
$+\{(1+{\alpha_{5}})(1-\omega)(1+\alpha_{8}^{-1})+{\kappa_{1}}\}\mathbb{E}\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}.$
where $\alpha_{5},\alpha_{8}>0$ and $\kappa_{1}$$\kappa_{3}$} are provided in Appendix 6.2.
Proof. See Appendix 6.2 for the proof.
Lemma 4
Suppose Assumptions 45 hold, one has
$\mathbb{E}$ $\|{\eta(\beta{\mathbf{m}_{t-1}}+{\mathbf{g}_{t-1}})}\|_{F}^{2}\leq\frac{{16{\eta^{2}}{\beta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{(1-\beta)\gamma_{g}^{2}}}\sum\limits_{k=0}^{t-1}{{\beta^{t-1-k}}}{\chi_{k}}$
$+\frac{{16{\eta^{2}}{d^{2}}\sigma_{2}^{2}L_{{f_{1}}}^{2}}}{{\gamma_{g}^{2}}}{\chi_{t-1}}+2{\eta^{2}}{\kappa_{4}}(1+\frac{{{\beta^{2}}}}{{{{(1-\beta)}^{2}}}}),$ (18)
where $\kappa_{4}$ can be found in the proof of Lemma 4.
Proof. See Appendix 6.2 for the proof.
Now, we show the convergence of CMSPZO algorithm under general nonconvex settings.
Theorem 1
Suppose Assumptions 15 hold. For the total number of iterations $T>{\tilde{\varepsilon}_{8}}$, each agent runs Algorithm 1 with
${m_{2}}<{\gamma_{x}}<{m_{1}},\penalty 10000\ \eta=(1-\beta)\sqrt{\frac{n}{T}},$
where $m_{1}$}, $m_{2}$} and $\tilde{\varepsilon}_{8}$} are provided in Appendix 6.4. Then,
$\frac{1}{T}$ $\sum\limits_{t=0}^{T-1}{\sum\limits_{i=1}^{n}{\mathbb{E}{{\|{{x_{i,t}}-{{\bar{x}}_{t}}}\|}^{2}}}}={\mathcal{O}}(\frac{1}{{T}})+{\mathcal{O}}(\frac{1}{{T}\gamma_{g}^{2}}),$ (19)
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}{{\|{\nabla{f}({{\bar{x}}_{t}})}\|}^{2}}$ $={\mathcal{O}}(\frac{1}{{\sqrt{nT}}})+{\mathcal{O}}(\frac{{\sqrt{n}}}{{\sqrt{T}}})+{\mathcal{O}}(\frac{{\sqrt{n}}}{{\sqrt{T}\gamma_{g}^{2}}})$
$+{\mathcal{O}}(\frac{{\sqrt{{n^{3}}}}}{{\sqrt{{T^{3}}}\gamma_{g}^{2}}})+{\mathcal{O}}(\frac{{\sqrt{{n^{3}}}}}{{\sqrt{{T^{3}}}\gamma_{g}^{4}}})+{\mathcal{O}}({{\gamma_{g}^{2}}}).$ (20)
Proof: The detailed proof is given in Appendix 6.4.
Remark 3
It is worth noting that the inherent noise of the single-point zero-order estimator and the use of constant step sizes, the RHS of (1) contains ${\mathcal{O}}({{\gamma_{g}^{2}}})$, which does not indicate a failure of Algorithm 1. Rather, it reflects the fact that under a noisy single-point estimator with constant step sizes, the algorithm can only converge to a neighborhood of a stationary point. The radius of this neighborhood is proportional to $\gamma_{g}$, $\sigma_{1}$, $\sigma_{2}$ and $\vartheta_{1}$.
From the right-hand side of (1), it follows immediately that sublinear convergence rate can be fulfilled if ${\gamma_{g}}={T^{-\frac{1}{8}}}$, which is presented in the following result.
Corollary 1
Under the same assumptions and parameters settings in Theorem 1, let ${\gamma_{g}}={T^{-\frac{1}{8}}}$ and $T>{\tilde{\varepsilon}_{9}}$, then,
$\frac{1}{T}$ $\sum\limits_{t=0}^{T-1}{\sum\limits_{i=1}^{n}{\mathbb{E}{{\|{{x_{i,t}}-{{\bar{x}}_{t}}}\|}^{2}}}}={\mathcal{O}}(\frac{1}{{T}})+{\mathcal{O}}(\frac{1}{{\sqrt[4]{{{T^{3}}}}}}),$ (21)
$\frac{1}{T}\sum\limits_{t=0}^{T-1}\mathbb{E}{{\|{\nabla{f}({{\bar{x}}_{t}})}\|}^{2}}=$ ${\mathcal{O}}(\frac{1}{{\sqrt{nT}}})+{\mathcal{O}}(\frac{{\sqrt{n}}}{{\sqrt{T}}})+{\mathcal{O}}(\frac{{\sqrt{n}}}{{\sqrt[4]{T}}})$
$+{\mathcal{O}}(\frac{{\sqrt{{n^{3}}}}}{{\sqrt[4]{{{T^{5}}}}}})+{\mathcal{O}}(\frac{{\sqrt{{n^{3}}}}}{T}),$ (22)
where $\tilde{\varepsilon}_{9}$} can be found in Appendix 6.5.
Proof: The detailed proof is given in Appendix 6.5.
4 Simulations
In this section, we provide the results of the simulations to prove the effectiveness of our algorithm.
First, the system consists of $n=6$ agents and dimension $d=30$ that exchange information through a connected undirected communication graph. And our theoretical results via numerical experiments on a nonconvex distributed binary classification task (Xie et al. (2024)):
$f_{i}(x,\xi_{i})={\hbar_{1}}(x)+{\hbar_{2}}(x),$ (23)
where ${\hbar_{1}}(x)=\sum_{j=1}^{m_{i}}\log(1+e^{-p_{ij}x^{\top}q_{ij}})$ denotes the logistic regression function, $m_{i}=200$ denotes the number of local observations per agent. $\xi_{i}=(p_{ij},q_{ij})$, $p_{ij}\in\mathbb{R}^{d}$ stands for the Gaussian features regenerated per iteration, $q_{ij}\in\{-1,1\}$ is labels. ${\hbar_{2}}(x)=\sum_{\iota=1}^{n}\frac{\varpi\kappa[x]_{\iota}^{2}}{1+\kappa[x]_{\iota}^{2}}$ represents a nonconvex regularizer with the regularization parameters $\varpi=0.001$ and $\kappa=1$. $x]_{\iota}$ is the $\iota$-th coordinate of $x\in\mathbb{R}^{d}$.
Moreover, we adopt the unbiased 2-bit quantizer in Wang et al. (2025), which satisfies Definition 1. The quantization operator $\mathcal{C}(\cdot)$ takes the form
$\mathcal{C}(x)=\frac{\|x\|_{\infty}}{2^{{k_{2}}-1}}\mathrm{sgn}(x)\circ\lfloor\frac{2^{{{k_{2}}}-1}|x|}{\|x\|_{\infty}}+{\varpi_{1}}\rfloor,$ (24)
where ${k_{2}}=2$, $\varpi_{1}$ is the random perturbation vector uniformly chosen from $0,1]^{d}$.
The performance metric
$P(T)=\min_{t\in T}\{\mathbb{E}_{\xi}\|\nabla f(\bar{x}_{t},\xi)\|^{2}+\tfrac{1}{n}\sum_{i=1}^{n}\|x_{i,t}-\bar{x}_{t}\|^{2}\}$
is evaluated in terms of both the iteration rounds and the communication bits among agents. The hyperparameters used in the experiments are summarized in Table 1.
We compare the Algorithm 1 against distributed stochastic gradient tracking method with a one-point gradient estimator (DSGT_\_1P) in Mhanna and Assaad (2023) and one-point zeroth-order stochastic compressed distributed primal-dual algorithm (DZOS_\_PDA).
Fig. 1 shows that CMSPZO converges slightly slower than DSGT_\_1P but still faster than DZOS_\_PDA. Furthermore, Fig. 2 illustrates that compared to other algorithms, CMSPZO converges to the same accuracy with fewer bits. This also demonstrates the efficiency of CMSPZO.
Table 1: Hyperparameters for Algorithm 1
Algorithm $\gamma_{x}$ $\eta$ $\beta$ $\gamma_{g}$
CMSPZO $0.3$ $0.02$ $0.9$ $0.99$
DZOS_\_PDA $0.3$ 0.02 - $0.99$
DSGT_\_1P - $0.02$ - $0.99$
Figure 1: Evolutions of $P(T)$ with respect to the number of iterations.
Figure 2: Evolutions of $P(T)$ with respect to the number of transmitted bits.
5 Conclusion
In this paper, we proposed the CMSPZO algorithm for stochastic distributed nonconvex optimization under the assumption that the gradient is not available and that only noisy single queries of the objective function are available at a time. We confirmed that the proposed algorithm achieves the convergence rate $\mathcal{O}(\frac{1}{\sqrt[4]{T}})$ with the exact solution under diminishing step sizes and can achieve the sublinear convergence rate ${\mathcal{O}}(\frac{1}{{\sqrt{T}}})$ towards a neighborhood of the stationary point under fixed step sizes. The simulation results imply that CMSPZO algorithm can achieve better performance than existing methods that do not incorporate momentum updates. Future directions include extending the algorithm to the one-point residual-feedback oracle or incorporating the Byzantine-robust aggregation rules.